### v1

# thanks: Talento-CM fellow

Abstract
With the advent of next-generation surveys, constraints on cosmological parameters are anticipated to become more stringent, particularly for the total neutrino mass. This study forecasts these constraints utilizing galaxy clusters from the Chinese Space Station Telescope (CSST). Employing Fisher matrix analysis, we derive the constraint $\sigma(M_{\nu})$ from cluster number counts, cluster power spectrum, and their combination. The investigation ranges from the standard cosmological model with neutrinos $\nu\Lambda$CDM to the inclusion of dynamic dark energy in the $\nu w_{0}w_{a}$ model, revealing a minor impact of dark energy on neutrino constraints. We examine the largest source of systematic uncertainty arising from the mass-observable relation and find that, with perfect knowledge of the scaling relation parameters, CSST clusters have the potential to enhance precision, tightening constraints to $0.034$ eV. We also study the effects of the maximum redshift $z_{max}$ and other uncertainties, including redshift, halo mass function, and bias. Furthermore, we emphasize the significance of accounting for the growth-induced scale-dependent bias (GISDB) effect, which is expected to have an impact of 1.5 to 2.2 times on the final constraint.
preprint: APS/123-QED
I introduction
In the standard model of particle physics, neutrinos are considered to be massless. However, a significant breakthrough in understanding neutrino properties occurred in 1998 when the Super-K atmospheric neutrino experiment in Japan confirmed the existence of oscillations [1], indicating that neutrinos can change from one flavor to another during their flight. This discovery provide evidence for non-zero neutrino masses, although they are extremely small.
Various oscillation experiments, including those involving solar, atmospheric, accelerator, and reactor neutrinos, have been conducted. These experiments can determine the differences in mass-squared values between the three different mass eigenstates of neutrinos. However, they cannot determine the absolute value of neutrino masses. As a result, neutrinos can be categorized into two possible scenarios: the normal hierarchy (NH), where the two lighter neutrinos are closer in mass, or the inverted hierarchy (IH), where the two heavier neutrinos are. The current neutrino oscillation experiment [2] measured:
$\Delta m_{21}^{2}\equiv m_{2}^{2}-m_{1}^{2}=7.5\times 10^{-5}eV^{2}$
$|\Delta m_{31}^{2}|\equiv|m_{3}^{2}-m_{1}^{2}|=2.5\times 10^{-3}eV^{2},$
in which the subscript {1,2,3} represents different mass eigenstate. So the oscillation experiment can constrain the total neutrino mass $M_{\nu}\equiv\sum_{\nu=1,2,3}m_{\nu}>0.06$ eV for NH($m_{3}>m_{2}>m_{1}$), and $M_{\nu}>0.10$ eV for IH($m_{2}>m_{1}>m_{3}$).
How to measure the absolute neutrino mass? There are three different approaches [3][4]: (i)neutrinoless double $\beta$-decay $(0\nu\beta\beta)$: the KamLAND-Zen experiment obtained $M_{\nu}<0.18-0.48$ eV at the 90% confidence level (C.L.) [5]; (ii)kinematic effects: the KATRIN tritium decay experiment obtained $M_{\nu}<1.1$ eV at 90% C.L. [6]. (iii)cosmology (see [7, 8] for reviews).
During the early time of the Universe, neutrinos exhibit relativistic behavior, leading to a delay at the time of matter-radiation equality. Consequently, this delay suppresses the growth of perturbations since the majority of growth takes place during the matter-dominated era. Additionally, due to their small mass, neutrinos tend to freely stream out of overdense regions, which is known as the free streaming effect . This effect smears out fluctuations on scales smaller than the free-streaming scale $k_{\text{fs}}$ [7]:
$k_{\text{fs}}=\frac{0.82}{\sqrt{1+z}}\left(\frac{m_{\nu}}{1eV}\right){{\,h\,{\rm Mpc}^{-1}}}.$ (1)
Considering the impacts of these factors on the expansion history of the Universe and the growth of cosmic structures, various techniques are employed to constrain the total neutrino mass, including the cosmic microwave background (CMB) primary anisotropies, CMB lensing, galaxy distribution, galaxy lensing, Lyman-$\alpha$ forest, 21 cm surveys, supernovae (see [9, 10] for reviews and their constraints). The current tightest constraint comes from combining observations of the Sloan Digital Sky Survey (SDSS) baryon acoustic oscillation (BAO) and redshift-space distortions (RSD), Planck CMB anisotropies, Pantheon Type Ia supernovae (SNe Ia), and Dark Energy Survey (DES) weak lensing and galaxy clustering measurements, resulting in an upper limit of $M_{\nu}<0.115$ eV at a 95% C.L. [11]. Many studies have been conducted to predict the constraints that can be achieved by next-generation surveys such as Large Synoptic Survey Telescope (LSST) or Vera C. Rubin Observatory [12], Euclid [13] and Square Kilometre Array (SKA) [14]. For instance, Yu et al. [15] find that a combination of LSST galaxies and CMB-S4 lensing can achieve constraints of $0.025$ eV, while Dey et al. [16] utilize 21-cm intensity mapping, galaxy clustering, cosmic shear observations from the SKA survey, and CMB primary anisotropies from CMB-S4 to obtain a constraint of $0.006$ eV. In this study, we focus on the probe of neutrino by using galaxy clusters [see 17, 18, for reviews] and provide a forecast for the Chinese Space Station Telescope (CSST), a next-generation survey[19, 20].
As the largest gravitationally bound structures in the Universe, galaxy clusters, although less numerous than galaxies, have more accurate theoretical descriptions and smaller observed redshift errors. The typical size of a cluster is about $1{{\,h^{-1}{\rm Mpc}}}$, which is comparable to the $k_{\text{fs}}$. Therefore clusters are sensitive to neutrino mass. Specifically, the two most studied observables of clusters are: cluster number counts $N(z)$ and cluster power spectrum $P_{c}(k)$. These two are complementary[21]: neutrinos not only suppress the growth rate of perturbations in the linear regime, leading to a decrease in $N(z)$, but also affect the fluctuation on small scales, consequently influencing the shape of $P_{c}(k)$.
CSST is a 2-meter space telescope positioned on the same orbit as the China Space Station. It will cover an area of 17,500 deg2 on the sky over a period of ten years. Compared to other large-field surveys, CSST excels in depth and image resolution[22]. Another unique advantage is its capability to conduct seamless spectroscopic observations within the same sky regions. With these advantages, CSST is expected to provide more extensive and higher-quality data. As a result, the improved data from CSST is anticipated to enable more accurate constraints on the total neutrino mass.
In this study, we aim to provide forecasts for constraints on the total neutrino mass $M_{\nu}$ using CSST galaxy clusters. The paper is organized as follows: In section II, we introduce our method, which is based on the Fisher matrix and incorporates neutrino prescriptions. Section III presents the main results of our analysis with different cosmological models and scenarios. In Section IV, we discuss factors that could impact our findings. Finally, we summarize and conclude our study in Section V.
Throughout this paper, the natural logarithm is denoted by $\ln$, while $\log$ denotes the decimal logarithm. The halo mass is defined as $M\equiv M_{200m}$, i.e. cluster mass enclosed within a spherical overdensity of 200 with respect to the mean matter density, unless otherwise specified.
II theoretical calculations
In this section, we present the theoretical framework that describes cluster number counts and cluster power spectrum, as well as their Fisher matrix. Particularly, we perform some modifications due to the presence of massive neutrinos.
II.1 cluster number counts
Galaxy clusters are regarded as dark matter halos. The halo mass function (hereafter, HMF) is defined to be the comoving number density of clusters, denoted as $n(M,z)$, as a function of cluster mass $\ln M$ and redshift $z$. The HMF is given by the equation:
$\frac{dn}{d\ln M}(M,z)=\frac{\rho_{m}}{M}\cdot f(\sigma)\left|\frac{d\ln\sigma}{d\ln M}\right|.$ (2)
Here, $f(\sigma)$ is a semi-analytical function calibrated against N-body simulations, which is related to the nonlinear collapse [23, 24, 25, 26, 27]. $\rho_{m}$ represents the matter density at the present day. The quantity $\sigma$ corresponds to the root mean square (rms) of the linear matter power spectrum $P_{m}$ and can be expressed as:
[…]
where, $W(kR)$ represents the Fourier transform of the top-hat window function with a radius $R$, defined as
$W(kR)=\frac{3[\sin(kR)-kR\cos(kR)]}{(kR)^{3}}.$ (4)
The relationship between the halo mass $M$ and the radius $R$ in the top-hat window function is given by:
$M=\frac{4\pi}{3}R^{3}\rho_{m}.$ (5)
The power spectrum $P_{m}$ in Eq.(3) is obtained from CLASS (Cosmic Linear Anisotropy Solving System) [28].
By integrating Eq.(2), we can determine the abundance of clusters within a specific mass and redshift range $(\Delta\ln M,\Delta z)$:
[…]
where, $\Delta\Omega$ represents the solid angle covered by the survey, and $dV(z)/dzd\Omega$ denotes the comoving volume per unit redshift and solid angle.
Considering the actual survey, there is always an error between the true and observed redshift $[z,z^{ob}]$, as well as for cluster mass $[M,M^{ob}]$.
For the observed redshift, we employ a Gaussian distribution $P(z^{ob}\mid z)$ with an expected value of $z$ and a variance of $\sigma^{2}_{z^{ob}\mid z}$. This distribution represents the probability that the observed photometric redshift $z^{ob}$ corresponds to the true redshift $z$:
$P(z^{ob}\mid z)=\frac{1}{\sqrt{2\pi}\sigma_{z^{ob}\mid z}}\exp\left[-\frac{(z^{ob}-z)^{2}}{2\sigma^{2}_{z^{ob}\mid z}}\right],$ (7)
$\sigma_{z^{ob}\mid z}(z)=\sigma_{z}(1+z),$ (8)
where the dispersion of photometric redshift $\sigma_{z^{ob}\mid z}(z)$ generally increases with redshift in the form of $1+z$ [29], originating from poorer redshift estimates for faint galaxies at higher redshifts.
For the observed cluster mass, following [30] [31], we assume that the probability distribution function of the observed mass $M^{ob}$ for clusters with a fixed mass $M$ and redshift $z$ is given by a log-normal distribution with an expected value $\ln M^{ob}|\ln M,z\right>$, and a variance $\sigma^{2}_{\ln M^{ob}\mid\ln M,z}$:
$P(\ln M^{ob}\mid\ln M,z)$ $=\frac{1}{\sqrt{2\pi}\sigma_{\ln M^{ob}|\ln M,z}}$
×$\times\exp\left[-\frac{x^{2}(\ln M^{ob},\ln M,z)}{2\sigma_{\ln M^{ob}\mid\ln M,z}^{2}}\right],$ (9)
where $x(\ln M^{ob},\ln M,z)\equiv\ln M^{ob}-\left<\ln M^{ob}|\ln M,z\right>$ represents the deviation of the observed mass from its mean value. The mean value $\ln M^{ob}|\ln M,z\right>$ is given by:
$\ln M^{ob}|\ln M,z\right>=\ln M+\ln M_{\text{bias}},$ (10)
where $\ln M_{\text{bias}}$ is the mass bias, parameterized as [31]:
$\ln M_{\text{bias}}=B_{M,0}+\alpha\ln(1+z).$ (11)
[…]
$\ln\lambda^{ob}|M,z\right>$ $=\ln A_{\lambda}+B_{\lambda}\ln(\frac{M}{3\times 10^{14}M_{\odot}h^{-1}})$
$+C_{\lambda}\ln(\frac{1+z}{1+0.45}),$ (12)
Parameters $A_{\lambda},B_{\lambda}$, $C_{\lambda}$ and $D_{\lambda}$ here are also nuisance parameters. To estimate $\sigma_{\ln\lambda^{ob}}^{2}$, we use $D_{\lambda}$ in $\sigma_{\ln\lambda}$ to replace the constant term and add an additional term to represent the projection effects. For simplicity, we replace the Poisson term $1/\left<\lambda^{ob}(M)\right>$ with $1/\exp\left<\ln\lambda^{ob}(M)\right>$:
$\sigma_{\ln\lambda^{ob}}^{2}=D_{\lambda}^{2}+\frac{1}{\exp\left<\ln\lambda^{ob}(M)\right>}.$ (13)
Finally, the scatter in the observed mass can be expressed as:
$\sigma_{\ln M^{ob}}^{2}=\frac{\sigma_{\ln\lambda^{ob}}^{2}}{B_{\lambda}}+\kappa(1+z)^{2}$ (14)
where the first term represents the scatter in the true mass $\sigma_{\ln M}$, derived from Eq.(II.1) and Eq.(13), and the second term accounts for an additional redshift-dependent projection effect, which is associated with the photometric redshift error and the algorithm employed by cluster finders [30]. The redshift coefficient $\kappa$ is a nuisance parameter.
It should be noted that, Costanzi et al. [36] use $M_{500c}$, defined as the cluster mass enclosed within a spherical overdensity of 500 with respect to the critical matter density. So, when applying Eq.(II.1), we convert mass from $M_{200m}$ to $M_{500c}$ assuming that the halo density profile is described by the Navarro-Frenk-White (NFW) model [38] with a concentration parameter from [39].
Let $\Delta\ln M^{ob}_{i}$ denotes the observed mass bin $i$, and $\Delta z_{j}^{ob}$ denotes the observed redshift bin $j$. Finally, the expectation value of the number of galaxy clusters $N_{ij}$ incorporating observed effects is given by:
$N_{ij}$ $\equiv N(\Delta\ln M^{ob}_{i},\Delta z_{j}^{ob})$
[…]
$\int d\ln M\frac{dn}{d\ln M}(M,z)P(\ln M^{ob}|\ln M,z).$
II.2 cluster power spectrum
The cluster power spectrum can be calculated from the linear matter power spectrum $P_{m}$ [40]:
$P_{c}(k^{\text{fid}}_{\perp},k^{\text{fid}}_{\|},z^{ob})$ $=b_{\text{eff}}^{2}(z)P_{m}(k^{\text{true}},z)$
$\cdot\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|},z)\mathcal{AP}(z)\mathcal{Z}(z),$ (16)
with a bias $b_{\text{eff}}(z)$ and incorporating several effects, the linear redshift-space distortion (RSD) effect $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|})$, Alcock-Paczynski (AP) effect $\mathcal{AP}(z)$, and redshift uncertainty $\mathcal{Z}(z)$. The wave numbers here are distinguished as fiducial $k^{\text{fid}}$ and true $k^{\text{true}}$, associated with the AP effect, while all wave numbers discussed in subsection II.1 are true $k=k^{\text{true}}$. Each term is detailed below.
The effective bias $b_{\text{eff}}(z)$ is averaged over all cluster bias $b(M,z)$ at redshift $z$:
$b_{\text{eff}}(z)=\left[\int d\ln M^{ob}\int d\ln M\frac{dn(M,z)}{d\ln M}P(\ln M^{ob}|\ln M,z)\right]^{-1}$
$\int d\ln M^{ob}\int d\ln M\frac{dn(M,z)}{d\ln M}P(\ln M^{ob}|\ln M,z)\cdot b(M,z),$ (17)
where, the mass observation effect $P(\ln M^{ob}|\ln M,z)$ has also been considered.
The RSD effect refers to the distortion of the observed clustering pattern of galaxies in redshift space due to their line-of-sight velocities, thus is closely related to the line-of-sight wave numbers $k^{\text{true}}_{\|}$:
$\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|},z)=\left[1+\beta(z)\left(\frac{k^{\text{true}}_{\|}}{k^{\text{true}}}\right)^{2}\right]^{2},$ (18)
in which, the redshift-distortion parameter is:
$\beta(z)\equiv\frac{1}{b_{\text{eff}}(z)}\frac{d\ln D_{\text{grow}}(z)}{d\ln a},$ (19)
where $D_{\text{grow}}$ is the linear growth rate, and $a$ is the expansion factor normalized to unity today.
The AP effect arises from assuming a fiducial cosmology to infer the position of a galaxy from its observed redshift and angle. If the ’true’ background cosmology differs from the fiducial cosmology, the inferred comoving radial and transverse distances are distorted with respect to the true ones. We fix the fiducial wave numbers $k^{\text{fid}}$, then compute the true wave numbers $k^{\text{true}}$:
$k_{\perp}^{\text{true}}$ $=k^{\text{fid}}_{\perp}\cdot D_{A}^{\text{fid}}(z)/D_{A}(z),$
$k_{\|}^{\text{true}}$ $=k^{\text{fid}}_{\|}\cdot H(z)/H^{\text{fid}}(z),$ (20)
and the AP effect term:
$\mathcal{AP}(z)=\left(\frac{H(z)}{H^{\text{fid}}(z)}\right)\left(\frac{D_{A}^{\text{fid}}(z)}{D_{A}(z)}\right)^{2},$ (21)
[…]
The treatment for redshift uncertainty in the cluster power spectrum is different from that in the number counts Eq.(7). The cluster number distribution is ’smeared’ along the radial direction, damping the power spectrum by [41]:
$\mathcal{Z}(z)=\exp\left\{-\left[\frac{k^{\text{true}}_{\|}}{H(z)}\sigma_{z^{ob}\mid z}(z)\right]^{2}\right\},$ (22)
where $\sigma_{z^{ob}\mid z}(z)$ is the photometric error varying with redshift Eq.(8).
II.3 neutrino prescription
Constraining cosmological parameters through galaxy clusters requires an accurate theory that involves nonlinear scales. Press-Schechter (PS) theory[23] firstly provides a universal formulation for describing the HMF Eq.(2): $f(\nu)=\sqrt{\frac{2}{\pi}}\nu e^{-\nu^{2}/2}$. Here, $\nu\equiv\delta_{\text{crit}}/\sigma(M,z)$ is peak-height, a measure of the rarity of halo, and $\delta_{\text{crit}}=1.686$ is the critical density for halo collapse, weakly dependent on the underlying cosmological model. Universality means that cosmological and redshift dependence enter the equation only through the peak height $\nu$, without affecting functional form [23, 24, 42]. However, subsequent N-body simulations have revealed deviations from universality in redshift [26, 43, 44] and in cosmological model [44].
Regarding cosmological models incorporating massive neutrinos, [45, 46, 47] have discovered that the impact of neutrinos on universality can be minimized by replacing the total matter density $\rho_{m}=\rho_{cdm+b+\nu}$ in Eq.(2) and Eq.(5) with the cold dark matter + baryon density $\rho_{cdm+b}$, [48, 49, 50], and replacing the total matter power spectrum $P_{m}=P_{cdm+b+\nu}$ in Eq.(3) with the cold dark matter + baryon power spectrum $P_{cdm+b}$. The physical reason behind this is that the clustering of neutrinos around the CDM overdensity is minimal and is unlikely to impact nonlinear collapse [51, 52].
By employing this CDM prescription, [47] shows that for simulations with massive neutrinos the HMF is well reproduced by the fitting formula of Tinker et al. [26], which we utilize in this paper.
According to the peak-background split (PBS) approach [53], the bias between the halo and the matter overdensities $b(M,z)$ in Eq.(17) can be derived from the HMF. Therefore, the prescription remains the same, i.e. replacing $P_{m}$ in Eq.(3) with $P_{cdm+b}$, which will recover the universality of bias on large scales [45, 46]. We follow [45] and continue to use the fitting formula proposed by Tinker et al. [27].
However, this prescription only partially recovers the universality of bias. The inclusion of massive neutrinos introduces scale dependence in bias due to two effects, both of which have been observed in N-body simulations: (i) the fact that halos trace CDM fluctuations rather than total (CDM + neutrino) matter density fluctuations [45, 46] and (ii) the scale-dependent growth of CDM density perturbations [54, 55] The former can also be eliminated by the CDM prescription, i.e. replacing $P_{m}$ in Eq.(II.2) with $P_{cdm+b}$. While, the latter, called growth-induced scale-dependent bias (GISDB) effect [56] requires further consideration. To incorporate the GISDB effect, we employ a fitting formula from Eq.(49) of [57], which is based on the spherical collapse model and the PBS approach [58, 57]. This method has been shown to be consistent with the N-body simulation [54]. Importantly, the scale-dependent bias obtained in [57] is nearly independent of mass or redshift:
$\frac{b^{L}(k)}{b^{L}(k_{\text{ref}})}$ $=[1+\Delta_{\Lambda CDM}\tanh(\alpha k/k_{\text{eq}})]$
$\cdot\left[1+\frac{\Delta_{L}}{2}\left(\tanh\left[\frac{\log(q)}{\Delta_{q}}\right]+1\right)\right]$
$\equiv R(k),$ (23)
where $\Delta_{\Lambda CDM}=4.8\times 10^{-3}$,$\alpha=4$, $\Delta_{q}=1.6$, $\Delta_{L}=0.55f_{\nu}$ with $f_{\nu}\equiv\Omega_{\nu}/\Omega_{m}$ representing the neutrino fraction , $q=5k/k_{\text{fs}}$ related to the free streaming scale Eq.(1). Here, $k_{\text{eq}}=0.015{{\,h\,{\rm Mpc}^{-1}}}$ represents the scale of matter-radiation equality, $k_{\text{ref}}$ is a reference wave number on larger scales. The superscript $L$ denotes Lagrangian bias. Since we use $P_{cdm+b}$ to calculate the cluster power spectrum, the Eulerian bias is given by $b=1+b^{L}$. Finally, we can transform the scale-independent effective bias $b_{\text{eff}}(z)=b_{\text{eff}}(k_{\text{ref}},z)$ from Eq.(17) into a scale-dependent effective bias:
$b_{\text{eff}}(z)\rightarrow b_{\text{eff}}(k,z)=[b_{\text{eff}}(k_{\text{ref}},z)-1]\cdot R(k)+1.$ (24)
Mistakenly ignoring these two neutrino-induced scale dependences in the bias, will have an impact on cosmological parameter inference in galaxy survey [59, 60, 61, 56]. In this paper, we focus on the latter effect, i.e. GISDB effect in cluster survey.
II.4 fisher matrix
We employ the Fisher matrix to assess the constraint capabilities of clusters. It is defined as the second derivatives of the logarithm of the likelihood function with respect to the parameters $(p_{\mu},p_{\nu})$:
$\mathcal{F}_{\mu\nu}\equiv-\left<\frac{\partial^{2}\ln\mathcal{L}}{\partial p_{\mu}\partial p_{\nu}}\right>,$ (25)
where $\mathcal{L}$ represents the likelihood function. Crudely speaking, the inverse of the Fisher matrix, denoted as $\mathcal{F}^{-1}$, can be considered as the best possible covariance matrix for the measurement errors on the parameters.
Assuming the observed cluster number count $D$ as a Poisson distribution with an expectation and variance $N$ [62], the logarithm of the likelihood function $\mathcal{L}(\bm{D})$ can be expressed as:
$\ln\mathcal{L}(\bm{D})=\sum_{ij}D_{ij}\ln N_{ij}-N_{ij}-\ln D_{ij}!,$ (26)
where the subscripts ${i,j}$ represent observed mass and redshift bins, $N_{ij}$ denotes the expected number count in the fiducial model in bin $(i,j)$ Eq.(II.1). The Fisher matrix for the cluster number count $N$, denoted as $\mathcal{F^{N}_{\mu\nu}}$, is then given by:
$\mathcal{F}_{\mu\nu}^{N}=\sum_{i,j}\frac{\partial N_{ij}}{\partial p_{\mu}}\frac{\partial N_{ij}}{\partial p_{\nu}}\frac{1}{N_{ij}}.$ (27)
For the observed cluster power spectrum, assuming it follows a Gaussian distribution with a mean value $P_{c}$ Eq.(II.2) and a fractional variance $\sigma_{P}^{2}/P_{c}^{2}=2/\left(V_{k}V_{\text{eff}}\right)$ [40, 21], the Fisher matrix, denoted as $\mathcal{F}_{\mu\nu}^{P}$, can be expressed as:
$\mathcal{F}_{\mu\nu}^{P}=\sum_{i,j}\frac{\partial\ln P_{c,ij}}{\partial p_{\mu}}\frac{\partial\ln P_{c,ij}}{\partial p_{\nu}}\frac{\left(V_{k}V_{\mathrm{eff}}\right)_{ij}}{2},$ (28)
where the subscripts $\{i,j\}$ represent two-dimensional $k^{\text{fid}}$-space cells and observed redshift bins, respectively.
$V_{\text{eff}}$ represents the effective volume probed by the survey [63, 64]:
[…]
$\left[\frac{P_{c}(k^{\text{fid}}_{\perp},k^{\text{fid}}_{\|},z^{ob})}{\mathcal{AP}(z^{ob})\bar{n}^{-1}(z^{ob})+P_{c}(k^{\text{fid}}_{\perp},k^{\text{fid}}_{\|},z^{ob})}\right]^{2}.$ (30)
The impact of redshift uncertainty on the volume factor is negligible. For simplicity in calculations, we have disregarded it, i.e. $dV^{ob}/(dz^{ob}d\Omega)=dV/(dzd\Omega)$ without a integration over $dzP(z^{ob}|z)$. However, integration over $dzP(z^{ob}|z)$ is necessary only for $\bar{n}$, the expected average number density (similar to Eq.(II.1)):
[…]
and $V_{k}$ represents the cylindrical volume factor in $k$-space:
$V_{k}(k^{\text{fid}}_{\perp},k^{\text{fid}}_{\|})=\frac{2\pi\Delta(k^{\text{fid}}_{\perp})^{2}\Delta k^{\text{fid}}_{\|}}{(2\pi)^{3}}.$ (32)
II.5 parameters
Table 1: Parameters and fiducial values.
Parameter Description Fiducial values
Cosmological parameters
$\Omega_{b}h^{2}$ Baryon density 0.02242
$\Omega_{m}h^{2}$ Matter density 111including massive neutrinos 0.14240
$\Omega_{\nu}h^{2}$ Massive neutrinos density 0.0006442
$\Omega_{DE}$ DE density 0.6889
$w_{0}$ DE EoS parameter -1
$w_{a}$ DE EoS parameter 0
$n_{s}$ Spectral index 0.9665
$\sigma_{8}$ Normalization of perturbations 0.8102
Scaling relation parameters
$A_{\lambda}$ Amplitude in $\ln\lambda^{ob}\right>$ 222mass-richness relation Eq.(II.1) 79.8
$B_{\lambda}$ Mass dependence in $\ln\lambda^{ob}\right>$ 2 0.93
$C_{\lambda}$ Redshift dependence in $\ln\lambda^{ob}\right>$ 2 -0.49
$D_{\lambda}$ Intrinsic scatter in $\sigma_{\ln\lambda^{ob}}$ 333mass-richness relation Eq.(13) 0.217
$\kappa$ Projection effect term in $\sigma_{\ln M^{ob}}$ 444projection effect Eq.(14) 0.01
$B_{M0}$ Constant term in $M_{\text{bias}}$ 555mass bias Eq.(11) 0
$\alpha$ Redshift dependence in $M_{\text{bias}}$ 5 0
Derived parameter
$M_{\nu}(eV)$ Total neutrino mass 0.06
The basic cosmological model we consider is an extension of the spatially flat $\Lambda$CDM model by adding one massive neutrino, label as $\nu\Lambda$CDM model. We also consider $\nu w_{0}w_{a}$ model, incorporating dynamic dark energy (DE) with an equation of state (EoS) represented by $w(a)=w_{0}+(1-a)w_{a}$ [65, 66]. The set of cosmological parameters encompasses:
$\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{\nu}h^{2},\Omega_{DE},w_{0},w_{a},n_{s},\sigma_{8}\},$ (33)
which denotes the baryon density, matter density, neutrino density, DE density, DE Eos parameters, scalar power-law index and present-day normalization of the primordial power spectrum, respectively. The fiducial values for $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$ are following [67], while $\Omega_{\nu}h^{2}=0.0006442$ corresponding to $M_{\nu}=0.06$ eV through:
$\Omega_{\nu}h^{2}=\frac{M_{\nu}}{93.14eV},$ (34)
and fiducial values for DE Eos parameters are set as $w_{0}=-1,w_{a}=0$.
Additionally, there exists a set of nuisance parameters associated with the scaling relation (SR):
$\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda},\kappa,B_{M0},\alpha\},$ (35)
where the first four parameters $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda}\}$ pertain to the mass-richness relation in Eq.(II.1) and Eq.(13), and their fiducial vales follow Table 2 of [36]: { 79.8, 0.93, -0.49, 0.217}. The projection effect term in Eq.(14) $\kappa=0.1^{2}$ due to a possible contamination rate of about 10% at low redshift for massive clusters for CSST [30, 68]. The mass bias parameters in Eq.(11) are specified as $B_{M0}=0$ and $\alpha=0$.
All parameters and their fiducial values are summarized in Table 1.
For CSST, the solid angle is $\Delta\Omega=5.33$, corresponding to the wide survey area of $17500$ deg2. The cluster redshift uncertainty in Eq.(8) is assumed to be $\sigma_{z}=0.001$, considering that the photometric redshift uncertainty for most galaxies is expected to be around $0.02$ for CSST. The incorporation of additional slitless grating spectroscopic data will leads to a smaller one, and the redshift uncertainty of each cluster as a whole would be reduced after averaging the redshifts of the constituent galaxies. Thus, we adopt an optimistic value of $0.001$. The observed redshift is assumed to be up to $z_{\max}=1.5$, using either redshift-based method [69, 70] or color-based method [71]. The observed mass threshold is $M_{200m}\geq 0.836\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}$, converted from $M_{500c}\geq 0.7\times 10^{14}\,\rm{M_{\odot}}$ to guarantee a completeness of $\geq 90\%$ and a purity of $\geq 90\%$ for the CSST cluster sample [69].
The fiducial wave number ranges span from $k^{\text{fid}}_{\perp,\|}=0.005$ to $0.15$  Mpc-1, with a bin size of $\Delta k^{\text{fid}}_{\perp,\|}=0.005$  Mpc-1. The observed mass bin is $\Delta\ln M^{ob}[\,\rm{M_{\odot}}]=0.2$. For cluster number counts, the observed redshift bin size is $\Delta z^{ob}=0.05$, while for the cluster power spectrum, it is $\Delta z^{ob}=0.2$, ensuring an adequate number of clusters within each redshift bin.
By summing over the aforementioned bins, we obtain the Fisher matrices $\mathcal{F}_{\mu\nu}^{N}$ for the cluster number counts Eq.(27) and $\mathcal{F}_{\mu\nu}^{P}$ for power spectrum Eq.(28). Finally, we can calculate the $1\,\sigma$ error for a parameter $p_{\mu}$ as $\sigma(p_{\mu})=\sqrt{(\mathcal{F}^{-1})_{\mu\mu}}$, as shown in Table 2.
III results
Figure 1: Cluster number counts as a function of redshift $z$.
Table 2: Forecasted $1\,\sigma$ errors for all parameters described in subsection II.5. When a parameter is kept fixed, we use ’–’. The header labels the different cosmological models ($\nu\Lambda$CDM and $\nu w_{0}w_{a}$), scenarios (self-calibration and known SR), and probes (cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$). For complete descriptions, please refer to the main text.
model $\nu\Lambda$CDM $\nu w_{0}w_{a}$
scenario Self-calibration Known SR Self-calibration Known SR
probe $N$ $P$ $N+P$ $N$ $P$ $N+P$ $N$ $P$ $N+P$ $N$ $P$ $N+P$
$\Omega_{b}h^{2}$ 1.831 0.006 0.005 0.454 0.004 0.004 1.968 0.008 0.005 0.468 0.005 0.004
$\Omega_{m}h^{2}$ 5.741 0.026 0.022 1.379 0.018 0.015 6.200 0.039 0.023 1.412 0.022 0.015
$\Omega_{\nu}h^{2}$ 0.056 0.002 0.001 0.008 0.002 3.4e-04 0.059 0.003 0.001 0.012 0.002 3.6e-04
$\Omega_{DE}$ 0.122 0.014 0.008 0.004 0.008 0.001 0.161 0.028 0.010 0.017 0.017 0.004
$w_{0}$ 0.083 0.168 0.057 0.067 0.130 0.050
$w_{a}$ 1.158 0.664 0.225 0.463 0.483 0.191
$n_{s}$ 3.790 0.046 0.039 0.827 0.028 0.017 4.177 0.069 0.040 0.836 0.035 0.017
$\sigma_{8}$ 0.258 0.018 0.010 0.010 0.004 0.002 0.269 0.036 0.010 0.058 0.006 0.003
$A_{\lambda}$ 1.5e+02 1.1e+07 1.1e+01 - - - 2.0e+02 1.1e+07 1.2e+01 - - -
$B_{\lambda}$ 1.024 1.7e+05 0.203 - - - 1.356 1.7e+05 0.207 - - -
$C_{\lambda}$ 1.628 1.1e+05 0.328 - - - 2.294 1.0e+05 0.336 - - -
$D_{\lambda}$ 0.363 7.1e+04 0.074 - - - 0.415 7.0e+04 0.075 - - -
$\kappa$ 0.466 5.4e+03 0.129 - - - 0.539 5.3e+03 0.136 - - -
$B_{M0}$ 0.860 6.4e+01 0.029 - - - 1.006 6.3e+01 0.033 - - -
$\alpha$ 0.431 1.9e+03 0.045 - - - 0.434 1.9e+03 0.047 - - -
$M_{\nu}(eV)$ 5.215 0.201 0.117 0.706 0.194 0.031 5.458 0.293 0.127 1.095 0.223 0.034
In this section, we present the main results based on the CSST survey. The redshift distribution of CSST clusters is shown in Figure 1, with a total number count of 408,037. The predicted $1\,\sigma$ errors for all parameters, especially for the total neutrino mass $M_{\nu}$, are shown in Table 2, for the various probes, cosmological models, and scenarios. Specifically, these constraints is derived from the cluster number counts $N$, the cluster power spectrum $P$, and their combination $N+P$, in the two cosmological models: the $\nu\Lambda$CDM model and the $\nu w_{0}w_{a}$ model. To explore the impact of the main systematics in cluster cosmology, i.e. the uncertainty in the SR, we consider two scenarios: (1) the self-calibration scenario, where the SR parameters are jointly constrained with the cosmological parameters; and (2) the known SR scenario, where the SR parameters are fixed.
To express clearly, the constraints on $M_{\nu}$ obtained from $N$, $P$ and $N+P$ are denoted as $\sigma^{N}(M_{\nu})$, $\sigma^{P}(M_{\nu})$ and $\sigma^{N+P}(M_{\nu})$ respectively.
III.1 $\nu\Lambda$CDM model
Figure 2: Constraints on cosmological parameters in the $\nu\Lambda$CDM model, with marginalizing over the SR parameters if exit. All contours denote 68% C.L. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenario respectively.
We begin with the $\nu\Lambda$CDM model, referring to columns 2-7 in Table 2. The corresponding contour plot is shown in Fig. 2 for the cosmological parameters $\{M_{\nu},\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$.
It is evident that constraints on all cosmological parameters primarily arise from $P$, except for $\Omega_{DE}$ in the known SR scenario. This is because $N$ is only sensitive to the integration of the matter power spectrum, rather than its shape, resulting in weak constraints on $\Omega_{b}h^{2}$, $n_{s}$ and $M_{\nu}$ [40, 21, 72]. Consequently, the constraint on $M_{\nu}$ from $N$ primarily stems from the influence of neutrinos on the amplitude of the matter power spectrum, leading to a strong degeneracy between $M_{\nu}$ and $\sigma_{8}$. In summary, the constraint on $M_{\nu}$ from $N$ is very weak, only $\sigma^{N}(M_{\nu})=5.215\ (0.706)$ eV in the self-calibration (known SR) scenario.
The power spectrum $P$ serves as a more sensitive probe of the neutrino mass compared to $N$, providing a constraint of $\sigma^{P}(M_{\nu})=0.201\ (0.706)$ eV on its own. When combined with $N$, the improvement factor in the self-calibration (known SR) scenario is 1.72 (6.19).
The incorporation of SR information has a substantial impact on $N$, resulting in an enhancement of constraints on all cosmological parameters by a factor of at least four. Particularly, the constraint on $M_{\nu}$ experiences a significant strength by a factor of 7.4, shifting from $5.215$ to $0.706$ eV. In contrast, the influence of SR information on $P$ is minimal, which is reflected in large errors of SR parameters in the third column, since the SR parameters only affects $b_{\text{eff}}$. Consequently, the enhance in the constraint on $M_{\nu}$ is only 1.04 times, from $0.201$ to $0.194$ eV, when incorporating SR information. In the combined analysis, the inclusion of SR information strengthens the constraint from $0.117$ to $0.031$ eV, which corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
III.2 $\nu w_{0}w_{a}$ model
Figure 3: Similar to Fig. 2, but in $\nu w_{0}w_{a}$ model.
Next we extend the $\nu\Lambda$CDM model to the $\nu w_{0}w_{a}$ model in order to investigate the impact of the dynamic dark energy on the neutrino constraint, referring to columns 8-13 in Table 2. The corresponding contour plot is shown in Fig. 3 for the cosmological parameters $\{M_{\nu},w_{0},w_{a}\}$ with other parameters marginalized over.
It is expected that including the parameters $\{w_{0},w_{a}\}$ will loosen the constraint on $M_{\nu}$ since both dark energy and massive neutrinos influence the growth rate of structures in the Universe. As depicted in Fig. 3, there is a small positive correlation between $M_{\nu}$ and $w_{0}$, as well as a distinct negative correlation between $M_{\nu}$ and $w_{a}$, for both $N$ and $P$. However, the direction of degeneracy is slightly different between $N$ and $P$, resulting in the constraint from the combination $N+P$ being less affected by the presence of the dynamic dark energy [73].
Consider the known SR scenario as an example. Adding $\{w_{0},w_{a}\}$ results in a loosening of $\sigma(M_{\nu})$ by a factor of 1.552 and 1.1471 for $N$ and $P$, respectively. In contrast, the combination $N+P$ demonstrates a loosening factor of only 1.079. The same holds true for the self-calibration scenario, where the loosening factor for the combination $N+P$ is 1.084.
In summary, the constraint on the total neutrino mass, $M_{\nu}$, is mainly contributed by the cluster power spectrum, $P$. The inclusion of dynamic dark energy has minimal impact, whereas the availability of perfect SR knowledge plays a significant role.
IV discussions
Table 3: Forecasted the $1\,\sigma$ marginalized error for the the total neutrino for different conditions. The 4th row represents the fiducial result, same as shown in Table 2. The subsequent rows demonstrate the results under specific conditions: partial SR knowledge (5th row), $z_{max}=2$ (6th row), $\sigma_{z}=0.01$ (7th row), HMF uncertainty (8th row), bias uncertainty (9th row), and without considering the GISDB effect (10th row). The symbol ’/’ denotes that the constraint is the same as the fiducial condition.
model $\nu\Lambda$CDM $\nu w_{0}w_{a}$
scenario Self-calibration Known SR Self-calibration Known SR
probe $N$ $P$ $N+P$ $N$ $P$ $N+P$ $N$ $P$ $N+P$ $N$ $P$ $N+P$
fiducial 5.215 0.201 0.117 0.706 0.194 0.031 5.458 0.293 0.127 1.095 0.223 0.034
partial 2.950 0.201 0.116 / / / 4.590 0.292 0.125 / / /
$z_{max}=2$ 3.072 0.196 0.104 0.642 0.190 0.030 3.329 0.287 0.118 0.909 0.218 0.033
$\sigma_{z}=0.01$ 5.266 0.514 0.188 0.706 0.364 0.065 5.509 0.653 0.241 1.098 0.475 0.073
HMF 5.238 0.201 0.127 0.830 0.199 0.066 6.524 0.293 0.235 2.945 0.224 0.135
bias / 0.201 0.161 / 0.200 0.091 / 0.293 0.249 / 0.290 0.119
w/o GISDB / 0.187 0.140 / 0.158 0.064 / 0.277 0.191 / 0.229 0.075
In this section, we analyze several systematic effects and survey strategies, as presented in Table 3, and conclude by comparing our forecast with results from other surveys at the end of the section.
IV.1 SR parameters
Figure 4: The constraint on the total neutrino mass $\sigma(M_{\nu})$ as we increase prior information on each SR parameter $\{B_{\lambda},D_{\lambda},\kappa,B_{M0},\alpha\}$ denoted with different colors. Each step represents the addition of a prior in a logarithmic space, from $10^{4}$ to $10^{-4}$, which is wide enough to cover the ’totally ignorant’ to ’totally known’ cases. Solid and dashed lines represent $\nu w_{0}w_{a}$ and $\nu\Lambda$CDM model respectively. The upper and the lower panel represents the constraint from $N$ and $N+P$ respectively.
In our analysis, we do not directly incorporate the mass observable in CSST, i.e. the optical richness, into our model. Although the mass-richness relation (MRR) has been calibrated in numerous works (e.g. [32, 33, 34, 35, 36]), it will be affected by different surveys and cluster finders [74, 75]. Therefore it is inappropriate to select a MRR calibrated in one specific survey to another survey. Instead, we employ a simple parameterization to account for our current lack of knowledge regarding the details of the mass-observable relations and their potential dependence on redshift and mass. However, our approach is not entirely devoid of information. Rather than treating $\sigma_{\ln M}$ as a nuisance parameter, we relate it with $\sigma_{\ln\lambda}$, and assume $\sigma_{\ln\lambda}$ comprises a combination of a constant intrinsic scatter and a Poisson-like term. In our main analysis, we have selected a MRR to model the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$. However, in actual surveys, we will utilize a specific MRR to replace Eq.(10), and there will be no need for an additional MRR to account for the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$. Consequently, we fix parameters $\{A_{\lambda},B_{\lambda},C_{\lambda}\}$ in Eq.(II.1), while retaining $B_{\lambda}$ as a nuisance parameter in Eq.(14) to mimic this scenario, which we refer to as the partial self-calibration scenario.
The 5th row in Table 3 shows the result. Similar to the previous section, SR parameters have a greater impact on $N$ than on $P$. The partial self-calibration scenario improve the constraint on $M_{\nu}$ from $N$ by a factor of 1.8 (1.2) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. However, the improvement is minimal for the constraint from $P$ or $N+P$.
We proceed to investigate the effect of the remaining SR parameters, $\{B_{\lambda},D_{\lambda},\kappa,B_{M0},\alpha\}$, on $\sigma(M_{\nu})$ by adding the prior information, as depicted in Fig. 4. The improvement trend is similar for both $\nu\Lambda$CDM and $\nu w_{0}w_{a}$ model. The upper panel illustrates that $\sigma^{N}(M_{\nu})$ is primarily influenced by $B_{M0}$ and $\alpha$, which determine the total number counts. On the other hand, the lower panel demonstrates that $\sigma^{N+P}(M_{\nu})$ is more sensitive to the scatter parameters $B_{\lambda},D_{\lambda}$ and $\kappa$, which suggests we’d better select an observable with smaller scatter when using the combination probe $N+P$ to constrain the total neutrino mass.
Recently, we calibrate the intrinsic mass-richness relation of clusters in THE THREE HUNDRED hydrodynamic simulations [76]. Our findings indicate that employing a skewed Gaussian distribution to model the richness at a fixed halo mass results in a simpler and smaller scatter compared to the conventional use of a log-normal distribution. Once this scaling relation is validated with the CSST mock data, we intend to incorporate it into future cosmological analyses.
IV.2 $z_{max}$
Figure 5: The impact of $z_{max}$ on $\sigma(M_{\nu})$. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenario respectively. The upper panel shows the $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model. The lower panel shows the ratio of $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ to $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model.
In our fiducial analysis, we assume a maximum redshift of $z_{\max}=1.5$ for cluster detection with CSST. However, by improving the cluster selection algorithm or jointing analysis with other surveys like Euclid [31], it may be possible to extend the maximum redshift to $z_{\max}=2$ [30]. The 6th row in Table 3 shows the result with $z_{max}=2$. Compared to the fiducial result, the total number counts is increased to 434534, and the constraint from $N$ has been improved by a factor of $\sim$1.7 (1.1) in the self-calibration (known SR) scenario for both two cosmological models. However, the improvement for the constraint from $P$ is very limited, with a factor of only $\sim$1.02 across all cases. Since the constraint on $M_{\nu}$ primarily comes from $P$, the improvement in $\sigma^{N+P}(M_{\nu})$ is not significant. In the self-calibration (known SR) scenario, the improvement is $\sim$1.1 (1.03) for both cosmological models.
We also investigate a wide range of $z_{max}$, as depicted in Fig. 5. The upper panel shows the $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model. It is evident that $\sigma^{N}(M_{\nu})$ decreases significantly with the increase of $z_{max}$. Conversely, $\sigma^{P}(M_{\nu})$ remains relatively constant once $z_{max}$ reaches $\sim$1.4. This constancy arises due to the exponential decrease in the cluster number density $\bar{n}$ Eq.(31) with increasing redshift, resulting in a very small effective volume $V_{\text{eff}}$ Eq.(II.4) at high redshifts. Consequently, the constraint from the combination $N+P$, i.e. $\sigma^{N+P}(M_{\nu})$, exhibits a moderate decrease with increasing $z_{\text{max}}$.
The lower panel shows the ratio of $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ to $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model. In the known SR scenario, the ratio from $N$ is significantly larger than 1 until $z_{max}$ reaches $\sim$2. This is because at low redshifts, dark energy can counterbalance the impact of neutrinos on the background evolution. However, at high redshifts, the influence of dark energy becomes secondary, refer to Fig.4 in [77]. On the other hand, the ratio from $P$ remains nearly constant once $z_{max}$ reaches $\sim$1.4, which is still due to the limit of $\bar{n}$.
IV.3 redshift uncertainty
Figure 6: The impact of the redshift uncertainty $\sigma_{z}$ on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model. Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. Solid and dotted lines represent self-calibration and known SR scenario respectively.
The redshift uncertainty $\sigma_{z}$ is optimistically assumed to be 0.001 in our fiducial analysis. This assumption is based on the availability of spectroscopic redshifts for all clusters observed by the CSST. However, in cases where this assumption does not hold, we consider a more conservative estimate of 0.01 for $\sigma_{z}$. The constraint result using this conservative estimate is presented in the 7th row of Table 3. We also extend $\sigma_{z}$ up to 0.03 in Fig. 5, which is the photometric redshift error for Euclid at high redshift [72].
The redshift uncertainty $\sigma_{z}$ has a minimal impact on $\sigma^{N}(M_{\nu})$ due to the redshift bin size. Even for the largest redshift uncertainty $\sim$0.03, it is still smaller than the redshift bin $\Delta z=0.05$ we choose. However, the influence of $\sigma_{z}$ on $\sigma^{P}(M_{\nu})$ is noteworthy. This is attributed to the damping effect caused by the redshift uncertainty, which leads to the partial erasure of information in $P$. It is observed that the behavior of $\sigma^{N+P}(M_{\nu})$ aligns with that of $\sigma^{P}(M_{\nu})$. Quantitatively, the ratio between $\sigma^{N+P}(M_{\nu})$ with $\sigma_{z}=0.01$ and $\sigma_{z}=0.001$ is $\sim$2. These findings underscore the significance of precise redshift measurements in effectively constraining the total neutrino mass.
IV.4 HMF uncertainty
Although the CDM prescription we use in subsection II.3 has alleviate the impact of neutrinos on the universality of HMF, the HMF of Tinker we employ is accurate at the $\sim$5level [26]. Additionally, baryonic effects introduce an uncertainty of $\sim$10to the HMF [78, 79]. To account for these uncertainties, we incorporate two additional parameters $\{s,q\}$ to characterize the deviation of the Tinker’s HMF from the true HMF, following [80] and [81]:
$\frac{dn}{d\ln M}=\left(\frac{dn}{d\ln M}\right)_{\text{Tinker}}\left[s\log\left(\frac{M}{10^{13.8}{{\,h^{-1}{\rm Mpc}}}}\right)+q\right].$ (36)
The fiducial values are $s=0$ and $q=1$, degenerating the HMF to the Tinker’s.
The 8th row in Table 3 presents the constraints with additional HMF uncertainties. Taking the $\nu w_{0}w_{a}$ model as an example, compared to $\sigma(M_{\nu})$ in the fiducial condition, the constraints from $N$ have been relaxed by a factor of 1.2 (2.7) in the self-calibration (known SR) scenario, while the constraints from $P$ remain largely unchanged. However for the constraints from $N+P$, the relaxation becomes more significant by a factor of 1.85 (3.97). Similar conclusions hold for the $\nu\Lambda$CDM model. As a result, the accuracy of the HMF plays a crucial role in constraining $M_{\nu}$ using $N+P$. It is worth mentioning that current and future emulators, based on simulations with massive neutrinos, have the potential to determine the HMF with greater precision, bypassing the reliance on a universal formula [82].
IV.5 bias uncertainty
Galaxy clusters are biased tracers of the CDM field. Fortunately, compared to galaxies, we have theory to describe the bias $b(M,z)$ in Eq.(17) for clusters, and the formula can be calibrated through simulations, similar to the HMF. Therefore, the bias of Tinker that we utilize also exhibits uncertainties. Moreover, in real surveys, extra bias will be introduced by selection effects, primarily due to projection effects and orientation biases. We incorporate two parameters $\{b_{s0},b_{s1}\}$ to model these uncertainties, following [83, 84]:
$b(M,z)=b_{\text{Tinker}}(M,z)b_{s0}\left(\frac{M}{5\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}}\right)^{b_{s1}},$ (37)
with fiducial values $b_{s0}=1,b_{s1}=0$.
The constraint on $M_{\nu}$ with the bias uncertainty is presented in the 9 th row in the Table 3. For $\sigma^{P}(M_{\nu})$, in the self-calibration scenario, the effect of bias uncertainty is overshadowed by the effect of SR parameters. However, in the known SR scenario, the inclusion of bias uncertainty relaxes the constraint by a factor of $\sim$1.03 (1.3) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. On the other hand, for $\sigma^{N+P}(M_{\nu})$, even in the self-calibration scenario, the effect of bias uncertainty is significant with a factor of $\sim$1.38(1.96). In the known SR scenario, it is 2.94 (3.5) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model. These findings highlight the greater influence of bias uncertainty compared to that of the HMF. It emphasizes the need for careful consideration of bias uncertainties in future analyses.
IV.6 GISDB
Figure 7: Bias. The impact of the GISDB effect on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model. Different colors (blue and black) represent different probes: cluster power spectrum $P$ and the combination $N+P$. Solid and hollow dots represent results with or without GISDB effect. Each column represents the self-calibration scenario (SC), the known SR scenario (SR), SC with the bias uncertainty and SR with the bias uncertainty, respectively.
In this section, we investigate the impact of the GISDB effect, which has been omitted in previous studies (e.g. [72, 31, 85]). The constraint without GISDB is presented in the 10th row of Table 3.
GISDB effect compensates for the suppression of the power spectrum caused by neutrinos at small scales.As a result, we anticipate that the constraint obtained from $P$ will be relaxed when GISDB is considered. However, there is an exception, $\sigma^{P}(M_{\nu})$ changing from 0.229 to 0.223. The reason for this exception is that when using $P$ to constrain $M_{\nu}$, we utilize information not only from the shape of the power spectrum but also from its amplitude, predominantly through the effective bias $b_{\text{eff}}$. By adding bias uncertainties, as discussed in the previous subsection, the amplitude information is blurred, leaving only the shape information.
Taking the $\nu w_{0}w_{a}$ model as an example, according to Fig. 7, when bias uncertainties are added, $\sigma(M_{\nu})$ with GISDB is consistently larger than without GISDB, regardless of whether the constraint is derived from $P$ or $N+P$. However, when the parameters $\{b_{s0},b_{s1}\}$ are fixed, the results change. This is because the degeneracy between $M_{\nu}$ and $\{b_{s0},b_{s1}\}$ differs when GISDB is considered or not. Especially for $\sigma^{N+P}(M_{\nu})$, considering GISDB will strengthen the constraint by a factor of $\sim$1.5 (2.2) in the self-calibration (known SR) scenario.
IV.7 Comparison with other cluster surveys
In this subsection, we compare our results for CSST with others for LSST, Euclid, and CMB-S4. All the constraints discussed here are derived in the $\nu w_{0}w_{a}$ model, except for LSST, which is in the $\nu w_{0}$ model.
The first work using cluster number counts and cluster power spectrum to constrain the total neutrino mass was performed by Wang et al. for LSST, a shear-limited weak lensing survey with a sky area of 18,000 deg2 and a redshift range of $z=[0.1,1.4]$. They forecast $\sigma^{P}(M_{\nu})=0.71$ eV and $\sigma^{N+P}(M_{\nu})=0.42$ eV, with an improvement factor of $\sim$1.34 when adding $N$. In the self-calibration scenario, we obtain $\sigma^{P}(M_{\nu})=0.293$ eV and $\sigma^{N+P}(M_{\nu})=0.127$ eV, with a improvement factor of $\sim$2.31. The main difference between our results and those of Wang et al. lies in the improvement factors of 1.34 and 2.31, respectively. This disparity can be attributed to the fact that the total cluster number counts forecasted by Wang et al. is approximately $2\times 10^{5}$, which is lower than our estimate of $4\times 10^{5}$.
Next, we compare our results with those obtained by Sartoris et al. for Euclid, an optical survey with a sky area of 15,000 deg2 and a redshift range of $z=[0.2,2]$ [31]. They use a selection threshold of $N_{500,c}/\sigma_{\text{field}}$, a ratio of $N_{500,c}$ the number of cluster galaxies contained within a sphere of certain radius and $\sigma_{\text{field}}$ the rms of the field galaxy counts. They forecast a total cluster number counts of $2\times 10^{6}$ and $2\times 10^{5}$ for $N_{500,c}/\sigma_{\text{field}}=3$ and 5 respectively. Our $4\times 10^{5}$ lies between them. For the former selection $N_{500,c}/\sigma_{\text{field}}=3$, they forecast $\sigma^{N+P}(M_{\nu})=0.140$ eV in the self-calibration scenario. When in the known SR scenario, it becomes 0.121 with an improvement factor of $\sim$1.16. Our main result is $\sigma^{N+P}(M_{\nu})=0.127$ eV in self-calibration, slightly smaller than $0.140$ eV. However, perfect knowledge of SR parameters bring our constraint greater improvement of $\sim$3.7 to $\sigma^{N+P}(M_{\nu})=0.034$ eV. When extending $z_{max}$ from 1.5 to 2, our results become smaller $\sigma^{N+P}(M_{\nu})=0.118\ (0.033)$ eV in self-calibration (known SR) scenario. It is worth noting that if we do not consider the GISDB effect as they do, our result becomes $\sigma^{N+P}(M_{\nu})=0.191\ (0.075)$ eV, larger than their $0.140$ eV.
Finally, we compare our results with those obtained by Cromer et al. for CMB-S4, a Sunyaev-Zel’dovich (SZ) survey with a sky area of 10,000 deg2 and a redshift range of $z=[0.1,1.9]$ [85]. They forecast a total cluster number counts of $\sim 10^{5}$. By combining Planck primary CMB and extra mass calibration from CMB lensing, they forecast $\sigma^{N+P}(M_{\nu})\sim 0.1$ eV, which falls between our results of $0.127$ eV and $0.034$ eV in the self-calibration and known SR scenarios, respectively. It is worth noting that they find the improvement on $\sigma(M_{\nu})$ when adding $P$ is very small $\sim$0.5%. This may be due to the fact that they have already incorporated prior information from the primary CMB and additional mass calibration, resulting in the $N$ alone having a certain constraining ability. Furthermore, their redshift uncertainty is reported to be 0.01, which has a significant impact on the constraining ability of $P$. They also consider the bias uncertainty and find it has less than 1% effect on $\sigma(M_{\nu})$, mainly because $P$ has a minimal contribution on the constraint.
V conclusions
In this work, we forecast constraints on the total neutrino mass from CSST galaxy clusters. Specifically, we employ Fisher matrix analysis to derive the constraint $\sigma(M_{\nu})$ from the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$. In addition to considering the most basic cosmological extension, the standard cosmology plus neutrinos, $\nu\Lambda$CDM model, we also explore the inclusion of dynamic dark energy, $\nu w_{0}w_{a}$ model. Since the largest source of systematics in galaxy cluster cosmology arises from the mass-observable relation, we examine cases where the SR parameters are completely unknown and perfectly known as self-calibration scenario and known SR scenario. The main results are presented in Table 3. The results between $\nu\Lambda CDM$ and $\nu w_{0}w_{a}$ model are similar. Unless otherwise specified, the following results refer to $\nu w_{0}w_{a}$ model:
  • CSST is expected to detect 408037 clusters with the mass threshold $M_{200m}\geq 0.836\times 10^{14}M_{\odot}/h$ and a survey area of 17,500 deg2 in the redshift range of $z=[0,1.5]$ , as depicted in Fig. 1.
  • Probe $N$ provides a weak constraint on $M_{\nu}$, $\sigma^{N}(M_{\nu})=5.458\ (1.095)$ eV in the self-calibration (known SR) scenario. The constraint is mainly contributed by $P$, with $\sigma^{P}(M_{\nu})=0.293\ (0.223)$ eV. However, when $N$ is added, the constraint can be improved by a factor of 2.31 (6.56), resulting in $\sigma^{N+P}(M_{\nu})=$ 0.127 (0.034) eV.
  • Having knowledge of the SR parameters has a significant impact on $\sigma^{N}(M_{\nu})$ and consequently affects $\sigma^{N+P}(M_{\nu})$. While partial knowledge leads to a small improvement in $\sigma^{N+P}(M_{\nu})$ from $0.127$ eV to $0.125$ eV, having perfect knowledge results in a substantial improvement to $0.034$ eV. This corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
  • Increasing $z_{max}$ from 1.5 to 2.0 will detect more clusters, increasing the count from $4\times 10^{5}$ to $4.3\times 10^{5}$. While this change has a very small impact on $\sigma^{P}(M_{\nu})$, improving it by a factor of 1.02 (1.02) in the self-calibration (known SR) scenario, it significantly strengthens $\sigma^{N}(M_{\nu})$ by a factor of 1.64 (1.20). As a result, it also strengthens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.08 (1.03), leading to $\sigma^{N+P}(M_{\nu})=0.118(0.033)$ eV.
  • Among the uncertainties we considered, the redshift uncertainty $\sigma_{z}$ has the largest impact on $\sigma^{P}(M_{\nu})$, relaxing the constraint by a factor of 2.23 (2.13) when increasing $\sigma_{z}$ from 0.001 to 0.01. Consequently, this leads to a relaxation in $\sigma^{N+P}(M_{\nu})$ by a factor of 1.90 (2.15) in the self-calibration (known SR) scenario.
  • We also examine the impact of uncertainties in the HMF and bias. The HMF uncertainties primarily affect $\sigma^{N}(M_{\nu})$, while the bias uncertainties mainly affect $\sigma^{P}(M_{\nu})$. However, the effects of both uncertainties on $\sigma^{N+P}(M_{\nu})$ are comparable, leading to a relaxation of the constraints by a factor of approximately 2 (4) in the self-calibration (known SR) scenario
  • Finally, we investigate the impact of the GISDB effect, which has been omitted in the previous work. This effect compensates for the suppression of the power spectrum caused by neutrinos on small scales, thereby relaxing the constraint on neutrinos. However, taking into account the degeneracy between $M_{\nu}$ and bias, as well as the addition of the probe $N$, our results demonstrate that considering this effect actually tightens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.50 (2.21) in the self-calibration (known SR) scenario.
In summary, CSST clusters are expected to provide a constraint on $M_{\nu}$ of $0.127$ eV. With perfect knowledge of the SR parameters, this constraint can be strengthened to $0.034$ eV, enabling a $2\sigma$ detection of the minimum neutrino mass of 0.06 eV. Recently, Lin et al. [86] presented a forecast using mock data from the CSST photometric galaxy clustering and cosmic shear surveys, which yielded an upper limit for $M_{\nu}$ of approximately $0.23$ eV at a 68% C.L. By combining clusters and galaxies, it is expected that CSST will significantly strengthen the constraint on the total neutrino mass. Moreover, the emergence of new probes such as voids [87] and Minkowski functions [88, 89] has opened up additional avenues for studying neutrino masses. By combining these innovative probes with traditional methods, it is anticipated that in the near future it will be possible to accurately constrain neutrino masses and distinguish between different mass hierarchies.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China Grants No. 12173036, 11773024, 11653002, 11421303, by the National Key R&\&&D Program of China Grant No. 2021YFC2203100, by the China Manned Space Project Grant No. CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No. WK3440000004 and WK3440000005, and by the CAS Interdisciplinary Innovation Team.

### v2

@@
 With the advent of next-generation surveys, constraints on cosmological parameters are anticipated to become more stringent, particularly for the total neutrino mass.
-This study forecasts these constraints utilizing galaxy clusters from the Chinese Space Station Telescope (CSST).
-Employing Fisher matrix analysis, we derive the constraint $\sigma(M_{\nu})$ from cluster number counts, cluster power spectrum, and their combination.
-The investigation ranges from the standard cosmological model with neutrinos $\nu\Lambda$CDM to the inclusion of dynamic dark energy in the $\nu w_{0}w_{a}$ model, revealing a minor impact of dark energy on neutrino constraints.
-We examine the largest source of systematic uncertainty arising from the mass-observable relation and find that, with perfect knowledge of the scaling relation parameters, CSST clusters have the potential to enhance precision, tightening constraints to $0.034$ eV.
-We also study the effects of the maximum redshift $z_{max}$ and other uncertainties, including redshift, halo mass function, and bias.
-Furthermore, we emphasize the significance of accounting for the growth-induced scale-dependent bias (GISDB) effect, which is expected to have an impact of 1.5 to 2.2 times on the final constraint.
+This study forecasts such constraints utilizing galaxy clusters from the Chinese Space Station Telescope (CSST).
+Employing Fisher matrix techniques, we derive the constraint $\sigma(M_{\nu})$ from cluster number counts, cluster power spectrum, and their combination.
+The investigation covers both the standard cosmological model with massive neutrinos $\nu\Lambda$CDM and the inclusion of dynamic dark energy in the $\nu w_{0}w_{a}$ model, revealing a minor impact of dark energy on neutrino mass constraints.
+We examine the largest source of systematic arising from the mass-observable relation uncertainties and find that, with perfect knowledge of the scaling relation parameters, CSST clusters have the potential to enhance precision, tightening constraints to $\sim 0.03$ eV.
+We also study the effects of the maximum redshift $z_{max}$ and other uncertainties, including those in redshift, halo mass function, and bias.
+Furthermore, we emphasize the significance of accounting for the growth-induced scale-dependent bias (GISDB) effect, which we find can tighten the final constraint by a factor of $1.2$ $2.2$.
 preprint: APS/123-QED
 I introduction
-In the standard model of particle physics, neutrinos are considered to be massless.
-However, a significant breakthrough in understanding neutrino properties occurred in 1998 when the Super-K atmospheric neutrino experiment in Japan confirmed the existence of oscillations [1], indicating that neutrinos can change from one flavor to another during their flight.
-This discovery provide evidence for non-zero neutrino masses, although they are extremely small.
-Various oscillation experiments, including those involving solar, atmospheric, accelerator, and reactor neutrinos, have been conducted.
-These experiments can determine the differences in mass-squared values between the three different mass eigenstates of neutrinos.
+In the standard model of particle physics, neutrinos are expected to be massless.
+However, a significant breakthrough in understanding neutrinos occurred in 1998 when the Super-Kamiokande atmospheric neutrino experiment in Japan confirmed neutrino oscillations [1], that is, neutrinos can change from one flavor to another during their propagation.
+This discovery provides evidence for non-zero neutrino masses, although small.
+Various oscillation experiments, including those involving solar, atmospheric, accelerator and reactor neutrinos, have been conducted.
+They can determine the differences in squared masses between the three different mass eigenstates of neutrinos.
 However, they cannot determine the absolute value of neutrino masses.
-As a result, neutrinos can be categorized into two possible scenarios: the normal hierarchy (NH), where the two lighter neutrinos are closer in mass, or the inverted hierarchy (IH), where the two heavier neutrinos are.
-The current neutrino oscillation experiment [2] measured:
+As a result, neutrinos can be organized into two possible scenarios: the normal hierarchy (NH), where the two lighter neutrinos are closer in mass, or the inverted hierarchy (IH), where the two heavier neutrinos are.
+The current neutrino oscillation experiments [2] measured:
 $\Delta m_{21}^{2}\equiv m_{2}^{2}-m_{1}^{2}=7.5\times 10^{-5}eV^{2}$
 $|\Delta m_{31}^{2}|\equiv|m_{3}^{2}-m_{1}^{2}|=2.5\times 10^{-3}eV^{2},$
-in which the subscript {1,2,3} represents different mass eigenstate.
-So the oscillation experiment can constrain the total neutrino mass $M_{\nu}\equiv\sum_{\nu=1,2,3}m_{\nu}>0.06$ eV for NH($m_{3}>m_{2}>m_{1}$), and $M_{\nu}>0.10$ eV for IH($m_{2}>m_{1}>m_{3}$).
+[…]
 How to measure the absolute neutrino mass?
-There are three different approaches [3][4]: (i)neutrinoless double $\beta$-decay $(0\nu\beta\beta)$: the KamLAND-Zen experiment obtained $M_{\nu}<0.18-0.48$ eV at the 90% confidence level (C.L.) [5]; (ii)kinematic effects: the KATRIN tritium decay experiment obtained $M_{\nu}<1.1$ eV at 90% C.L.
+There are three different approaches [3][4]: (i) neutrinoless double $\beta$-decay $(0\nu\beta\beta)$ searches: the KamLAND-Zen experiment obtained the lightest neutrino mass $m_{\text{lightest}}<0.18-0.48$ eV at the 90% confidence level (C.L.) [5]; (ii) kinematic measurements: the KATRIN tritium decay experiment obtained $M_{\nu}<1.1$ eV at 90% C.L.
 [6].
 (iii) cosmology (see [7, 8] for reviews).
-During the early time of the Universe, neutrinos exhibit relativistic behavior, leading to a delay at the time of matter-radiation equality.
-Consequently, this delay suppresses the growth of perturbations since the majority of growth takes place during the matter-dominated era.
-Additionally, due to their small mass, neutrinos tend to freely stream out of overdense regions, which is known as the free streaming effect .
-This effect smears out fluctuations on scales smaller than the free-streaming scale $k_{\text{fs}}$ [7]:
-$k_{\text{fs}}=\frac{0.82}{\sqrt{1+z}}\left(\frac{m_{\nu}}{1eV}\right){{\,h\,{\rm Mpc}^{-1}}}.$ (1)
-Considering the impacts of these factors on the expansion history of the Universe and the growth of cosmic structures, various techniques are employed to constrain the total neutrino mass, including the cosmic microwave background (CMB) primary anisotropies, CMB lensing, galaxy distribution, galaxy lensing, Lyman-$\alpha$ forest, 21 cm surveys, supernovae (see [9, 10] for reviews and their constraints).
-The current tightest constraint comes from combining observations of the Sloan Digital Sky Survey (SDSS) baryon acoustic oscillation (BAO) and redshift-space distortions (RSD), Planck CMB anisotropies, Pantheon Type Ia supernovae (SNe Ia), and Dark Energy Survey (DES) weak lensing and galaxy clustering measurements, resulting in an upper limit of $M_{\nu}<0.115$ eV at a 95% C.L.
-[11].
-Many studies have been conducted to predict the constraints that can be achieved by next-generation surveys such as Large Synoptic Survey Telescope (LSST) or Vera C.
-Rubin Observatory [12], Euclid [13] and Square Kilometre Array (SKA) [14].
+Fixing the matter density today, if neutrinos have mass, the non-relativistic component will decrease in the early Universe because neutrinos exhibit relativistic behavior at that time.
+This will lead to a delay in the time of matter-radiation equality.Consequently, this delay suppresses the growth of perturbations since the majority of growth takes place during the matter-dominated era.
+Additionally, due to their large thermal velocities, neutrinos tend to freely stream out of overdense regions, which is known as the free streaming effect.
+This effect tends to smear out fluctuations on scales smaller than the free-streaming scale $k_{\text{fs}}$ [7]:
+$k_{\text{fs}}=0.82\frac{\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}}{(1+z)^{2}}\left(\frac{m_{\nu}}{1eV}\right){{\,h\,{\rm Mpc}^{-1}}},$ (1)
+where $\Omega_{\Lambda}$ (resp. $\Omega_{m}$) is the cosmological constant (resp. matter) density fraction evaluated today.
+Considering the impacts of neutrinos on the expansion history of the Universe and the growth of cosmic structures, various probes are developed to constrain the total neutrino mass, including the cosmic microwave background (CMB) anisotropies, CMB lensing, galaxy clustering, galaxy lensing, Lyman-$\alpha$ forest, 21 cm intensity mapping (see [9, 10] for reviews).
+The current tightest constraint comes from combining baryon acoustic oscillation (BAO) data from the Dark Energy Spectroscopic Instrument (DESI), with CMB anisotropies and CMB lensing data from both Planck and Atacama Cosmology Telescope (ACT), resulting in an upper limit of $M_{\nu}<0.072(0.113)$ eV at a 95% C.L.
+for a $M_{\nu}>0(0.059)$ eV prior [11].
+Many studies have been conducted to forecast constraints that can be achieved by next-generation surveys such as the Vera Rubin Observatory Legacy Survey of Space and Time (LSST) [12], Nancy Grace Roman Space Telescope [13] and Square Kilometre Array (SKA)[14].
 For instance, Yu et al.
-[15] find that a combination of LSST galaxies and CMB-S4 lensing can achieve constraints of $0.025$ eV, while Dey et al.
-[16] utilize 21-cm intensity mapping, galaxy clustering, cosmic shear observations from the SKA survey, and CMB primary anisotropies from CMB-S4 to obtain a constraint of $0.006$ eV.
-In this study, we focus on the probe of neutrino by using galaxy clusters [see 17, 18, for reviews] and provide a forecast for the Chinese Space Station Telescope (CSST), a next-generation survey[19, 20].
-As the largest gravitationally bound structures in the Universe, galaxy clusters, although less numerous than galaxies, have more accurate theoretical descriptions and smaller observed redshift errors.
-The typical size of a cluster is about $1{{\,h^{-1}{\rm Mpc}}}$, which is comparable to the $k_{\text{fs}}$.
-Therefore clusters are sensitive to neutrino mass.
+[15] find that a combination of LSST galaxies and CMB-S4 lensing can achieve constraints on $M_{\nu}$ of $0.025$ eV.
+In this study, we focus on the probe of neutrino mass by using galaxy clusters [see 16, 17, for reviews] and provide a forecast for the Chinese Space Station Telescope (CSST), a next-generation space-based galaxy survey[18, 19].
+As the largest gravitationally bound structures in the Universe, galaxy clusters, although less abundant than galaxies, have more accurate theoretical descriptions and smaller observed redshift errors.
+Clusters are also sensitive to neutrino mass.
 Specifically, the two most studied observables of clusters are: cluster number counts $N(z)$ and cluster power spectrum $P_{c}(k)$.
-These two are complementary[21]: neutrinos not only suppress the growth rate of perturbations in the linear regime, leading to a decrease in $N(z)$, but also affect the fluctuation on small scales, consequently influencing the shape of $P_{c}(k)$.
-CSST is a 2-meter space telescope positioned on the same orbit as the China Space Station.
-It will cover an area of 17,500 deg2 on the sky over a period of ten years.
-Compared to other large-field surveys, CSST excels in depth and image resolution[22].
-Another unique advantage is its capability to conduct seamless spectroscopic observations within the same sky regions.
-With these advantages, CSST is expected to provide more extensive and higher-quality data.
-As a result, the improved data from CSST is anticipated to enable more accurate constraints on the total neutrino mass.
-In this study, we aim to provide forecasts for constraints on the total neutrino mass $M_{\nu}$ using CSST galaxy clusters.
-The paper is organized as follows: In section II, we introduce our method, which is based on the Fisher matrix and incorporates neutrino prescriptions.
-Section III presents the main results of our analysis with different cosmological models and scenarios.
-In Section IV, we discuss factors that could impact our findings.
-Finally, we summarize and conclude our study in Section V.
-Throughout this paper, the natural logarithm is denoted by $\ln$, while $\log$ denotes the decimal logarithm.
+These two are complementary[20]: massive neutrinos not only suppress the overall amplitude of matter fluctuations, leading to a decrease in $N(z)$, but also affect the fluctuations on scales smaller than the free-streaming scales, consequently influencing the shape of $P_{c}(k)$.
+CSST is a 2-meter space telescope operating in the same orbit as the China Manned Space Station.
+It will perform simultaneously a multi-band imaging and slitless spectroscopic survey with an area of 17,500 deg2 in about ten years.
+Compared to other large-field surveys, CSST excels in image quality, number of filter bands, and its unique coverage in the near ultraviolet etc [21].
+Hence, CSST is expected to be highly competitive.
+It will also be complementary to other large projects of its time, such as Euclid [22] and the Roman Space Telescope [13].
+As a result, we expect CSST can provide more accurate constraints on the total neutrino mass as well.
+In this work, we aim at forecasting constraints on the total neutrino mass $M_{\nu}$ by using CSST galaxy clusters, specifically the two cluster probes of cluster number counts and power spectrum.
+The rest of this paper is organized as follows: in section II, we introduce our methodology, which is based on the Fisher matrix technique, with careful incorporation of effects from massive neutrinos.
+Section III presents our main results for different models and assumptions.
+In Section IV, we discuss various effects that could impact our findings.
+Finally, we summarize our results and conclude in Section V.
+Throughout this paper, natural logarithm is denoted as $\ln$, while $\log$ stands for decimal logarithm.
 The halo mass is defined as $M\equiv M_{200m}$, i.e.
-cluster mass enclosed within a spherical overdensity of 200 with respect to the mean matter density, unless otherwise specified.
+cluster mass enclosed within a radius with an interior overdensity of 200 times the mean matter density, unless otherwise specified.
 II theoretical calculations
-In this section, we present the theoretical framework that describes cluster number counts and cluster power spectrum, as well as their Fisher matrix.
-Particularly, we perform some modifications due to the presence of massive neutrinos.
+In this section, we present our theoretical calculations for cluster number counts and cluster power spectrum, as well as their Fisher matrices.
+Particularly, we incorporate the effects of massive neutrinos on both observables.
 II.1 Cluster number counts
-Galaxy clusters are regarded as dark matter halos.
-The halo mass function (hereafter, HMF) is defined to be the comoving number density of clusters, denoted as $n(M,z)$, as a function of cluster mass $\ln M$ and redshift $z$.
-The HMF is given by the equation:
+Galaxy clusters can be taken as dark matter halos, with their comoving number density $n$ described by the halo mass function (hereafter, HMF), denoted as $dn/d\ln M(M,z)$, as a function of cluster mass $\ln M$ and redshift $z$:
 $\frac{dn}{d\ln M}(M,z)=\frac{\rho_{m}}{M}\cdot f(\sigma)\left|\frac{d\ln\sigma}{d\ln M}\right|.$ (2)
-Here, $f(\sigma)$ is a semi-analytical function calibrated against N-body simulations, which is related to the nonlinear collapse [23, 24, 25, 26, 27].
+Here, $f(\sigma)$ is a function calibrated against N-body simulations, which is related to the nonlinear collapse [23, 24, 25, 26, 27].
 $\rho_{m}$ represents the matter density at the present day.
-The quantity $\sigma$ corresponds to the root mean square (rms) of the linear matter power spectrum $P_{m}$ and can be expressed as:
+The quantity $\sigma(R,z)$ is the amplitude of the linear matter fluctuations at redshift $z$, smoothed by a top hat window function with scale $R$, and can be expressed as:
 […]
-where, $W(kR)$ represents the Fourier transform of the top-hat window function with a radius $R$, defined as
+where, $W(kR)$ represents the Fourier transform of the top-hat window function with a radius $R$, given by
 $W(kR)=\frac{3[\sin(kR)-kR\cos(kR)]}{(kR)^{3}}.$ (4)
-The relationship between the halo mass $M$ and the radius $R$ in the top-hat window function is given by:
+The relationship between the halo mass $M$ and the scale $R$ in the top-hat window function is given by:
 $M=\frac{4\pi}{3}R^{3}\rho_{m}.$ (5)
-The power spectrum $P_{m}$ in Eq.(3) is obtained from CLASS (Cosmic Linear Anisotropy Solving System) [28].
+The linear matter power spectrum $P_{m}$ in Eq.(3) is obtained from CLASS (Cosmic Linear Anisotropy Solving System) [28].
 By integrating Eq.(2), we can determine the abundance of clusters within a specific mass and redshift range $(\Delta\ln M,\Delta z)$:
 […]
-where, $\Delta\Omega$ represents the solid angle covered by the survey, and $dV(z)/dzd\Omega$ denotes the comoving volume per unit redshift and solid angle.
-Considering the actual survey, there is always an error between the true and observed redshift $[z,z^{ob}]$, as well as for cluster mass $[M,M^{ob}]$.
-For the observed redshift, we employ a Gaussian distribution $P(z^{ob}\mid z)$ with an expected value of $z$ and a variance of $\sigma^{2}_{z^{ob}\mid z}$.
-This distribution represents the probability that the observed photometric redshift $z^{ob}$ corresponds to the true redshift $z$:
+where, $\Delta\Omega$ represents the solid angle covered by the survey, and $dV(z)/dzd\Omega$ denotes the comoving volume per unit redshift and solid angle intervals.
+Considering an actual survey, there is always an error between the true and observed redshifts $[z,z^{ob}]$, as well as for cluster mass $[M,M^{ob}]$.
+For the observed redshift, we employ a Gaussian distribution for $z^{ob}$, $P(z^{ob}\mid z)$, with an expected value of $z$ and a variance of $\sigma^{2}_{z^{ob}\mid z}$:
 $P(z^{ob}\mid z)=\frac{1}{\sqrt{2\pi}\sigma_{z^{ob}\mid z}}\exp\left[-\frac{(z^{ob}-z)^{2}}{2\sigma^{2}_{z^{ob}\mid z}}\right],$ (7)
 $\sigma_{z^{ob}\mid z}(z)=\sigma_{z}(1+z),$ (8)
-where the dispersion of photometric redshift $\sigma_{z^{ob}\mid z}(z)$ generally increases with redshift in the form of $1+z$ [29], originating from poorer redshift estimates for faint galaxies at higher redshifts.
-For the observed cluster mass, following [30] [31], we assume that the probability distribution function of the observed mass $M^{ob}$ for clusters with a fixed mass $M$ and redshift $z$ is given by a log-normal distribution with an expected value $\ln M^{ob}|\ln M,z\right>$, and a variance $\sigma^{2}_{\ln M^{ob}\mid\ln M,z}$:
+where the dispersion of observed redshift $\sigma_{z^{ob}\mid z}(z)$ generally increases with redshift with a factor of $1+z$ [29], reflecting poorer redshift estimates for faint galaxies at higher redshifts.
+For the observed cluster mass, following [30] [31], we assume that the probability distribution function of the observed mass $M^{ob}$ for clusters with a fixed true mass $M$ and true redshift $z$ is given by a log-normal distribution with an expected value $\ln M^{ob}\right>(\ln M,z)$, and a variance $\sigma^{2}_{\ln M^{ob}\mid\ln M,z}$:
 $P(\ln M^{ob}\mid\ln M,z)$ $=\frac{1}{\sqrt{2\pi}\sigma_{\ln M^{ob}|\ln M,z}}$
 ×$\times\exp\left[-\frac{x^{2}(\ln M^{ob},\ln M,z)}{2\sigma_{\ln M^{ob}\mid\ln M,z}^{2}}\right],$ (9)
-where $x(\ln M^{ob},\ln M,z)\equiv\ln M^{ob}-\left<\ln M^{ob}|\ln M,z\right>$ represents the deviation of the observed mass from its mean value. The mean value $\ln M^{ob}|\ln M,z\right>$ is given by:
-$\ln M^{ob}|\ln M,z\right>=\ln M+\ln M_{\text{bias}},$ (10)
-where $\ln M_{\text{bias}}$ is the mass bias, parameterized as [31]:
+where $x(\ln M^{ob},\ln M,z)\equiv\ln M^{ob}-\left<\ln M^{ob}\right>(\ln M,z)$ represents the deviation of the observed mass from its mean value.
+The mean value $\ln M^{ob}\right>(\ln M,z)$ is given by:
+$\ln M^{ob}\right>(\ln M,z)=\ln M+\ln M_{\text{bias}},$ (10)
+where $\ln M_{\text{bias}}$ is the bias in the logarithm of mass, parameterized as [31]:
 $\ln M_{\text{bias}}=B_{M,0}+\alpha\ln(1+z).$ (11)
+The parameters $B_{M,0}$ and $\alpha$ here will be simultaneously constrained with cosmological parameters, often referred to as the self-calibration scheme.
 […]
-$\ln\lambda^{ob}|M,z\right>$ $=\ln A_{\lambda}+B_{\lambda}\ln(\frac{M}{3\times 10^{14}M_{\odot}h^{-1}})$
-$+C_{\lambda}\ln(\frac{1+z}{1+0.45}),$ (12)
-Parameters $A_{\lambda},B_{\lambda}$, $C_{\lambda}$ and $D_{\lambda}$ here are also nuisance parameters.
-To estimate $\sigma_{\ln\lambda^{ob}}^{2}$, we use $D_{\lambda}$ in $\sigma_{\ln\lambda}$ to replace the constant term and add an additional term to represent the projection effects.
-For simplicity, we replace the Poisson term $1/\left<\lambda^{ob}(M)\right>$ with $1/\exp\left<\ln\lambda^{ob}(M)\right>$:
-$\sigma_{\ln\lambda^{ob}}^{2}=D_{\lambda}^{2}+\frac{1}{\exp\left<\ln\lambda^{ob}(M)\right>}.$ (13)
-Finally, the scatter in the observed mass can be expressed as:
+$\ln\lambda^{ob}\right>(M,z)$ $=\ln A_{\lambda}+B_{\lambda}\ln(\frac{M}{3\times 10^{14}M_{\odot}h^{-1}})$
+$+C_{\lambda}\ln(\frac{1+z}{1+0.45}).$ (12)
+Without the projection effects, we would estimate $\sigma_{\ln\lambda^{ob}}^{2}$ also as the sum of a constant term, which we denote as $D_{\lambda}^{2}$ as well, and a Poisson term. That is,
+$\sigma_{\ln\lambda^{ob}}^{2}=D_{\lambda}^{2}+\frac{1}{\exp\left<\ln\lambda^{ob}(M)\right>},$ (13)
+which after scaled by the square of the scaling-relation slope $B_{\lambda}^{2}$, will lead to a scatter in the logarithm of mass.
+Finally, we model $\sigma_{\ln M^{ob}}^{2}$ as the direct sum of this scaled scatter and an extra term accounting for the projection effects:
 $\sigma_{\ln M^{ob}}^{2}=\frac{\sigma_{\ln\lambda^{ob}}^{2}}{B_{\lambda}}+\kappa(1+z)^{2}$ (14)
-where the first term represents the scatter in the true mass $\sigma_{\ln M}$, derived from Eq.(II.1) and Eq.(13), and the second term accounts for an additional redshift-dependent projection effect, which is associated with the photometric redshift error and the algorithm employed by cluster finders [30].
-The redshift coefficient $\kappa$ is a nuisance parameter.
+where the second term accounts for the redshift-dependent projection effect that is associated with redshift errors and the algorithm employed by the cluster finder [30].
 It should be noted that, Costanzi et al.
-[36] use $M_{500c}$, defined as the cluster mass enclosed within a spherical overdensity of 500 with respect to the critical matter density.
+[36] use $M_{500c}$, defined as the cluster mass enclosed within a radius with interior mean density of 500 times the critical density.
 So, when applying Eq.(II.1), we convert mass from $M_{200m}$ to $M_{500c}$ assuming that the halo density profile is described by the Navarro-Frenk-White (NFW) model [38] with a concentration parameter from [39].
-Let $\Delta\ln M^{ob}_{i}$ denotes the observed mass bin $i$, and $\Delta z_{j}^{ob}$ denotes the observed redshift bin $j$.
-Finally, the expectation value of the number of galaxy clusters $N_{ij}$ incorporating observed effects is given by:
+Let $\Delta\ln M^{ob}_{i}$ denote the $i$th observed mass bin, and $\Delta z_{j}^{ob}$ denote the $j$th observed redshift bin.
+Finally, the expectation value of the number of detected galaxy clusters $N_{ij}$ is given by:
 $N_{ij}$ $\equiv N(\Delta\ln M^{ob}_{i},\Delta z_{j}^{ob})$
@@
 II.2 Cluster power spectrum
-The cluster power spectrum can be calculated from the linear matter power spectrum $P_{m}$ [40]:
+The cluster power spectrum $P_{c}$ can be calculated from the matter power spectrum $P_{m}$ [40]:
 $P_{c}(k^{\text{fid}}_{\perp},k^{\text{fid}}_{\|},z^{ob})$ $=b_{\text{eff}}^{2}(z)P_{m}(k^{\text{true}},z)$
 $\cdot\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|},z)\mathcal{AP}(z)\mathcal{Z}(z),$ (16)
-with a bias $b_{\text{eff}}(z)$ and incorporating several effects, the linear redshift-space distortion (RSD) effect $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|})$, Alcock-Paczynski (AP) effect $\mathcal{AP}(z)$, and redshift uncertainty $\mathcal{Z}(z)$.
-The wave numbers here are distinguished as fiducial $k^{\text{fid}}$ and true $k^{\text{true}}$, associated with the AP effect, while all wave numbers discussed in subsection II.1 are true $k=k^{\text{true}}$.
+with a bias factor $b_{\text{eff}}(z)$ and several other factors, accounting for the redshift-space distortion (RSD) effect $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|})$, Alcock-Paczynski (AP) effect $\mathcal{AP}(z)$, and redshift uncertainty $\mathcal{Z}(z)$.
+The wave numbers here are distinguished as fiducial ones $k^{\text{fid}}$ and true ones $k^{\text{true}}$, due to the AP effect.
 Each term is detailed below.
-The effective bias $b_{\text{eff}}(z)$ is averaged over all cluster bias $b(M,z)$ at redshift $z$:
+The effective bias $b_{\text{eff}}(z)$ is an average of individual cluster bias $b(M,z)$ at redshift $z$:
 $b_{\text{eff}}(z)=\left[\int d\ln M^{ob}\int d\ln M\frac{dn(M,z)}{d\ln M}P(\ln M^{ob}|\ln M,z)\right]^{-1}$
 $\int d\ln M^{ob}\int d\ln M\frac{dn(M,z)}{d\ln M}P(\ln M^{ob}|\ln M,z)\cdot b(M,z),$ (17)
-where, the mass observation effect $P(\ln M^{ob}|\ln M,z)$ has also been considered.
-The RSD effect refers to the distortion of the observed clustering pattern of galaxies in redshift space due to their line-of-sight velocities, thus is closely related to the line-of-sight wave numbers $k^{\text{true}}_{\|}$:
+where the mass measurement uncertainty has also been considered.
+The RSD effect refers to the distortion of the observed clustering pattern of clusters in redshift space due to their line-of-sight velocities, thus has an explicit dependence on the line-of-sight wave number $k^{\text{true}}_{\|}$:
 $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|},z)=\left[1+\beta(z)\left(\frac{k^{\text{true}}_{\|}}{k^{\text{true}}}\right)^{2}\right]^{2},$ (18)
@@
 $\beta(z)\equiv\frac{1}{b_{\text{eff}}(z)}\frac{d\ln D_{\text{grow}}(z)}{d\ln a},$ (19)
-where $D_{\text{grow}}$ is the linear growth rate, and $a$ is the expansion factor normalized to unity today.
-The AP effect arises from assuming a fiducial cosmology to infer the position of a galaxy from its observed redshift and angle.
+where $D_{\text{grow}}$ is the linear growth rate, and $a$ is the scale factor normalized to unity today.
+The AP effect arises from assuming a fiducial cosmology to infer the position of a cluster from its observed redshift and angular position.
 If the ’true’ background cosmology differs from the fiducial cosmology, the inferred comoving radial and transverse distances are distorted with respect to the true ones.
@@
 […]
-The treatment for redshift uncertainty in the cluster power spectrum is different from that in the number counts Eq.(7).
-The cluster number distribution is ’smeared’ along the radial direction, damping the power spectrum by [41]:
+The treatment to account for redshift uncertainty in the cluster power spectrum is different from that in the number counts.
+The cluster distribution is ’smeared’ along the radial direction, which damps the power spectrum by [41]:
 $\mathcal{Z}(z)=\exp\left\{-\left[\frac{k^{\text{true}}_{\|}}{H(z)}\sigma_{z^{ob}\mid z}(z)\right]^{2}\right\},$ (22)
-where $\sigma_{z^{ob}\mid z}(z)$ is the photometric error varying with redshift Eq.(8).
+where $\sigma_{z^{ob}\mid z}(z)$ is the cluster redshift uncertainty Eq.(8).
 II.3 Neutrino prescription
-Constraining cosmological parameters through galaxy clusters requires an accurate theory that involves nonlinear scales.
-Press-Schechter (PS) theory[23] firstly provides a universal formulation for describing the HMF Eq.(2): $f(\nu)=\sqrt{\frac{2}{\pi}}\nu e^{-\nu^{2}/2}$.
-Here, $\nu\equiv\delta_{\text{crit}}/\sigma(M,z)$ is peak-height, a measure of the rarity of halo, and $\delta_{\text{crit}}=1.686$ is the critical density for halo collapse, weakly dependent on the underlying cosmological model.
-Universality means that cosmological and redshift dependence enter the equation only through the peak height $\nu$, without affecting functional form [23, 24, 42].
-However, subsequent N-body simulations have revealed deviations from universality in redshift [26, 43, 44] and in cosmological model [44].
-Regarding cosmological models incorporating massive neutrinos, [45, 46, 47] have discovered that the impact of neutrinos on universality can be minimized by replacing the total matter density $\rho_{m}=\rho_{cdm+b+\nu}$ in Eq.(2) and Eq.(5) with the cold dark matter + baryon density $\rho_{cdm+b}$, [48, 49, 50], and replacing the total matter power spectrum $P_{m}=P_{cdm+b+\nu}$ in Eq.(3) with the cold dark matter + baryon power spectrum $P_{cdm+b}$.
+Galaxy clusters are non-linear objects that form at the peaks of the initial density field.
+The Press-Schechter (PS) theory[23] firstly provides a universal formulation for describing the HMF Eq.(2): $f(\nu)=\sqrt{\frac{2}{\pi}}\nu e^{-\nu^{2}/2}$.
+Here, $\nu\equiv\delta_{\text{crit}}/\sigma(M,z)$ specifies the peak height, a measure of the rarity of halo, and $\delta_{\text{crit}}=1.686$ is the critical density for halo collapse, weakly dependent on the underlying cosmological model.
+Universality means that cosmological and redshift dependences enter the equation only through the peak height $\nu$, without affecting the functional form [23, 24, 42].
+However, accurate N-body simulations have revealed deviations from universality in redshift [26, 43, 44] and in cosmological model [44].
+Regarding cosmological models incorporating massive neutrinos, [45, 46, 47] find that the impact of massive neutrinos can be minimized by replacing the total matter density $\rho_{m}=\rho_{cdm+b+\nu}$ in Eq.(2) and Eq.(5) with the cold dark matter + baryon density $\rho_{cdm+b}$, [48, 49, 50], and replacing the total matter power spectrum $P_{m}=P_{cdm+b+\nu}$ in Eq.(3) with the cold dark matter + baryon power spectrum $P_{cdm+b}$.
 The physical reason behind this is that the clustering of neutrinos around the CDM overdensity is minimal and is unlikely to impact nonlinear collapse [51, 52].
-By employing this CDM prescription, [47] shows that for simulations with massive neutrinos the HMF is well reproduced by the fitting formula of Tinker et al. [26], which we utilize in this paper.
-According to the peak-background split (PBS) approach [53], the bias between the halo and the matter overdensities $b(M,z)$ in Eq.(17) can be derived from the HMF.
-Therefore, the prescription remains the same, i.e.
-replacing $P_{m}$ in Eq.(3) with $P_{cdm+b}$, which will recover the universality of bias on large scales [45, 46].
+By employing this CDM prescription, [47] shows that for simulations with massive neutrinos the HMF is well reproduced by the fitting formula of Tinker et al.
+[26], which we will utilize in this paper.
+According to the peak-background split (PBS) approach [53], the bias between the halo and matter overdensities $b(M,z)$ in Eq.(17) can be derived from the HMF.
+Therefore, the prescription with respect to the bias for cosmologies with massive neutrinos remains the same, i.e.
+replacing $P_{m}$ with $P_{cdm+b}$, and $\rho_{m}$ with $\rho_{cdm+b}$, which can recover the halo bias on large scales for simulations with massive neutrinos [45, 46].
 We follow [45] and continue to use the fitting formula proposed by Tinker et al.
-[27].
-However, this prescription only partially recovers the universality of bias.
-The inclusion of massive neutrinos introduces scale dependence in bias due to two effects, both of which have been observed in N-body simulations: (i) the fact that halos trace CDM fluctuations rather than total (CDM + neutrino) matter density fluctuations [45, 46] and (ii) the scale-dependent growth of CDM density perturbations [54, 55] The former can also be eliminated by the CDM prescription, i.e.
+[27] for halo bias.
+However, this prescription only partially accounts for the effects of massive neutrinos on the large-scale bias.
+Massive neutrinos also introduce scale dependence in the bias due to two effects, both of which have been observed in N-body simulations: (i) halos trace CDM fluctuations rather than total (CDM + neutrino) matter fluctuations [45, 46], and (ii) massive neutrinos introduce scale-dependence in the growth of CDM density perturbations [54, 55].
+The former can be eliminated by excluding neutrinos when defining bias, i.e.
 replacing $P_{m}$ in Eq.(II.2) with $P_{cdm+b}$.
-While, the latter, called growth-induced scale-dependent bias (GISDB) effect [56] requires further consideration.
+While, the latter, called the growth-induced scale-dependent bias (GISDB) [56], requires careful treatment.
 To incorporate the GISDB effect, we employ a fitting formula from Eq.(49) of [57], which is based on the spherical collapse model and the PBS approach [58, 57].
-This method has been shown to be consistent with the N-body simulation [54].
-Importantly, the scale-dependent bias obtained in [57] is nearly independent of mass or redshift:
+This method has been shown to be consistent with simulations [54].
+The scale-dependence obtained by [57] is almost independent of mass and redshift [^1] The scale-dependence does show weak dependence on redshift, which is neglected in our calculation.
+We agree a more accurate treatment would be using Relicfast directly rather than the fitting formula of Eq.
+(II.3).:
 $\frac{b^{L}(k)}{b^{L}(k_{\text{ref}})}$ $=[1+\Delta_{\Lambda CDM}\tanh(\alpha k/k_{\text{eq}})]$
@@
 $\equiv R(k),$ (23)
-where $\Delta_{\Lambda CDM}=4.8\times 10^{-3}$,$\alpha=4$, $\Delta_{q}=1.6$, $\Delta_{L}=0.55f_{\nu}$ with $f_{\nu}\equiv\Omega_{\nu}/\Omega_{m}$ representing the neutrino fraction , $q=5k/k_{\text{fs}}$ related to the free streaming scale Eq.(1).
-Here, $k_{\text{eq}}=0.015{{\,h\,{\rm Mpc}^{-1}}}$ represents the scale of matter-radiation equality, $k_{\text{ref}}$ is a reference wave number on larger scales.
+where $\Delta_{\Lambda CDM}=4.8\times 10^{-3}$, $\alpha=4$, $\Delta_{q}=1.6$, $\Delta_{L}=0.55f_{\nu}$ with $f_{\nu}\equiv\Omega_{\nu}/\Omega_{m}$ representing the neutrino fraction, $q=5k/k_{\text{fs}}$ (related to the free streaming scale Eq.(1)).
+Here, $k_{\text{eq}}$ represents the scale of matter-radiation equality [60] [^2] Here, for simplicity, we treat $k_{eq}$ as a constant., and $k_{\text{ref}}$ is a reference wave number on large scales, which we set to be $10^{-4}{{\,h^{-1}{\rm Mpc}}}$.
 The superscript $L$ denotes Lagrangian bias.
 Since we use $P_{cdm+b}$ to calculate the cluster power spectrum, the Eulerian bias is given by $b=1+b^{L}$.
-Finally, we can transform the scale-independent effective bias $b_{\text{eff}}(z)=b_{\text{eff}}(k_{\text{ref}},z)$ from Eq.(17) into a scale-dependent effective bias:
+Finally, we can replace the scale-independent effective bias $b_{\text{eff}}(z)=b_{\text{eff}}(k_{\text{ref}},z)$ in Eq.(17) by the scale-dependent effective bias:
 $b_{\text{eff}}(z)\rightarrow b_{\text{eff}}(k,z)=[b_{\text{eff}}(k_{\text{ref}},z)-1]\cdot R(k)+1.$ (24)
-Mistakenly ignoring these two neutrino-induced scale dependences in the bias, will have an impact on cosmological parameter inference in galaxy survey [59, 60, 61, 56].
+Ignoring these two neutrino-induced scale dependences in the bias, will have an impact on cosmological parameter inference in galaxy survey [62, 63, 64, 56].
 In this paper, we focus on the latter effect, i.e.
@@
 II.4 Fisher matrix
-We employ the Fisher matrix to assess the constraint capabilities of clusters.
-It is defined as the second derivatives of the logarithm of the likelihood function with respect to the parameters $(p_{\mu},p_{\nu})$:
+We employ the Fisher matrix techniques to assess the constraining power of clusters.
+It is defined as the ensemble average of the second derivatives of the logarithm of likelihood function with respect to the parameters $(p_{\mu},p_{\nu})$:
 $\mathcal{F}_{\mu\nu}\equiv-\left<\frac{\partial^{2}\ln\mathcal{L}}{\partial p_{\mu}\partial p_{\nu}}\right>,$ (25)
 where $\mathcal{L}$ represents the likelihood function.
-Crudely speaking, the inverse of the Fisher matrix, denoted as $\mathcal{F}^{-1}$, can be considered as the best possible covariance matrix for the measurement errors on the parameters.
-Assuming the observed cluster number count $D$ as a Poisson distribution with an expectation and variance $N$ [62], the logarithm of the likelihood function $\mathcal{L}(\bm{D})$ can be expressed as:
+The inverse of the Fisher matrix, denoted as $\mathcal{F}^{-1}$, provides the best possible covariance matrix for the measurement errors on the parameters.
+Assuming the observed cluster number count $D$ obeys a Poisson distribution with an expected value and variance $N$ [65], the logarithm of the likelihood function $\mathcal{L}(\bm{D})$ can be expressed as:
 $\ln\mathcal{L}(\bm{D})=\sum_{ij}D_{ij}\ln N_{ij}-N_{ij}-\ln D_{ij}!,$ (26)
@@
 $\mathcal{F}_{\mu\nu}^{P}=\sum_{i,j}\frac{\partial\ln P_{c,ij}}{\partial p_{\mu}}\frac{\partial\ln P_{c,ij}}{\partial p_{\nu}}\frac{\left(V_{k}V_{\mathrm{eff}}\right)_{ij}}{2},$ (28)
-where the subscripts $\{i,j\}$ represent two-dimensional $k^{\text{fid}}$-space cells and observed redshift bins, respectively.
+where the subscripts $\{i,j\}$ represent the two-dimensional $\vec{k}^{\text{fid}}$-space cells and the observed redshift bins, respectively.
 $V_{\text{eff}}$ represents the effective volume probed by the survey [66, 67]:
@@
 II.5 Parameters
-Table 1: Parameters and fiducial values.
+Table 1: Parameters and their fiducial values.
 Parameter Description Fiducial values
@@
 $\Omega_{b}h^{2}$ Baryon density 0.02242
-$\Omega_{m}h^{2}$ Matter density 111including massive neutrinos 0.14240
+$\Omega_{m}h^{2}$ Matter density 333including massive neutrinos 0.14240
 $\Omega_{\nu}h^{2}$ Massive neutrinos density 0.0006442
@@
 Scaling relation parameters
-$A_{\lambda}$ Amplitude in $\ln\lambda^{ob}\right>$ 222mass-richness relation Eq.(II.1) 79.8
-$B_{\lambda}$ Mass dependence in $\ln\lambda^{ob}\right>$ 2 0.93
-$C_{\lambda}$ Redshift dependence in $\ln\lambda^{ob}\right>$ 2 -0.49
-$D_{\lambda}$ Intrinsic scatter in $\sigma_{\ln\lambda^{ob}}$ 333mass-richness relation Eq.(13) 0.217
-$\kappa$ Projection effect term in $\sigma_{\ln M^{ob}}$ 444projection effect Eq.(14) 0.01
-$B_{M0}$ Constant term in $M_{\text{bias}}$ 555mass bias Eq.(11) 0
-$\alpha$ Redshift dependence in $M_{\text{bias}}$ 5 0
+$A_{\lambda}$ Amplitude in $\ln\lambda^{ob}\right>$ 444see Eq.(II.1) 79.8
+$B_{\lambda}$ Mass dependence in $\ln\lambda^{ob}\right>$ 4 0.93
+$C_{\lambda}$ Redshift dependence in $\ln\lambda^{ob}\right>$ 4 -0.49
+$D_{\lambda}$ Intrinsic scatter in $\sigma_{\ln\lambda^{ob}}$ 555see Eq.(13) 0.217
+$\kappa$ Projection effect parameter 666see Eq.(14) 0.01
+$B_{M0}$ Constant term in $M_{\text{bias}}$ 777see Eq.(11) 0
+$\alpha$ Redshift dependence in $M_{\text{bias}}$ 7 0
 Derived parameter
 $M_{\nu}(eV)$ Total neutrino mass 0.06
-The basic cosmological model we consider is an extension of the spatially flat $\Lambda$CDM model by adding one massive neutrino, label as $\nu\Lambda$CDM model.
+The basic cosmological model we consider is an extension of the flat $\Lambda$CDM model by adding one species of massive neutrino, labeled as $\nu\Lambda$CDM model, with the other two species assumed to be massless.
 We also consider $\nu w_{0}w_{a}$ model, incorporating dynamic dark energy (DE) with an equation of state (EoS) represented by $w(a)=w_{0}+(1-a)w_{a}$ [68, 69].
@@
 $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{\nu}h^{2},\Omega_{DE},w_{0},w_{a},n_{s},\sigma_{8}\},$ (33)
-which denotes the baryon density, matter density, neutrino density, DE density, DE Eos parameters, scalar power-law index and present-day normalization of the primordial power spectrum, respectively.
-The fiducial values for $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$ are following [67], while $\Omega_{\nu}h^{2}=0.0006442$ corresponding to $M_{\nu}=0.06$ eV through:
+which denotes the baryon density, matter density, neutrino density, DE density, DE Eos parameters, scalar power spectrum index and present-day normalization of the primordial power spectrum, respectively.
+The fiducial values for $\{\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$ follows [60], while $\Omega_{\nu}h^{2}=0.0006442$ corresponding to $M_{\nu}=0.06$ eV through [7]:
 $\Omega_{\nu}h^{2}=\frac{M_{\nu}}{93.14eV},$ (34)
 and fiducial values for DE Eos parameters are set as $w_{0}=-1,w_{a}=0$.
-Additionally, there exists a set of nuisance parameters associated with the scaling relation (SR):
+Additionally, there are a set of parameters associated with the scaling relation (SR) uncertainty:
 $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda},\kappa,B_{M0},\alpha\},$ (35)
-where the first four parameters $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda}\}$ pertain to the mass-richness relation in Eq.(II.1) and Eq.(13), and their fiducial vales follow Table 2 of [36]: { 79.8, 0.93, -0.49, 0.217}.
-The projection effect term in Eq.(14) $\kappa=0.1^{2}$ due to a possible contamination rate of about 10% at low redshift for massive clusters for CSST [30, 68].
+where the first four parameters $\{A_{\lambda},B_{\lambda},C_{\lambda},D_{\lambda}\}$ pertain to the mass-richness scaling in Eq.(II.1) and Eq.(13), and their fiducial values follow Table IV and Eq.(B1) of [36]: { 79.8, 0.93, -0.49, 0.217}.
+For the projection effect term in Eq.(14), we set $\kappa=0.1^{2}$ accounting for a possible contamination rate of about 10% at low redshift for massive clusters for CSST [30, 70].
 The mass bias parameters in Eq.(11) are specified as $B_{M0}=0$ and $\alpha=0$.
 All parameters and their fiducial values are summarized in Table 1.
-For CSST, the solid angle is $\Delta\Omega=5.33$, corresponding to the wide survey area of $17500$ deg2.
-The cluster redshift uncertainty in Eq.(8) is assumed to be $\sigma_{z}=0.001$, considering that the photometric redshift uncertainty for most galaxies is expected to be around $0.02$ for CSST.
-The incorporation of additional slitless grating spectroscopic data will leads to a smaller one, and the redshift uncertainty of each cluster as a whole would be reduced after averaging the redshifts of the constituent galaxies.
+For CSST, the solid angle is $\Delta\Omega=5.33$ sr, corresponding to the wide survey area of $17500$ deg2.
+The cluster redshift uncertainty in Eq.(8) is assumed to be $\sigma_{z}=0.001$ [71], considering that the photometric redshift uncertainty for galaxies is expected to be around $0.02$ for CSST [72].
+The incorporation of additional slitless grating spectroscopic data will leads to a smaller error, and the redshift uncertainty of each cluster as a whole would be reduced after averaging the redshifts of the member galaxies.
 Thus, we adopt an optimistic value of $0.001$.
@@
 The observed mass threshold is $M_{200m}\geq 0.836\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}$, converted from $M_{500c}\geq 0.7\times 10^{14}\,\rm{M_{\odot}}$ to guarantee a completeness of $\geq 90\%$ and a purity of $\geq 90\%$ for the CSST cluster sample [73].
-The fiducial wave number ranges span from $k^{\text{fid}}_{\perp,\|}=0.005$ to $0.15$  Mpc-1, with a bin size of $\Delta k^{\text{fid}}_{\perp,\|}=0.005$  Mpc-1.
-The observed mass bin is $\Delta\ln M^{ob}[\,\rm{M_{\odot}}]=0.2$.
-For cluster number counts, the observed redshift bin size is $\Delta z^{ob}=0.05$, while for the cluster power spectrum, it is $\Delta z^{ob}=0.2$, ensuring an adequate number of clusters within each redshift bin.
+The fiducial wave number for cluster power spectrum spans from $k^{\text{fid}}_{\perp,\|}=0.005$ to $0.15$  Mpc-1, with a bin size of $\Delta k^{\text{fid}}_{\perp,\|}=0.005$  Mpc-1.
+For cluster number counts, the observed mass bin size is $\Delta\ln M^{ob}[\,\rm{M_{\odot}}]$ = 0.2, and the observed redshift bin size is $\Delta z^{ob}=0.05$.
+While for the cluster power spectrum, we choose $\Delta z^{ob}=0.2$, ensuring adequate number of clusters within each redshift bin.
 By summing over the aforementioned bins, we obtain the Fisher matrices $\mathcal{F}_{\mu\nu}^{N}$ for the cluster number counts Eq.(27) and $\mathcal{F}_{\mu\nu}^{P}$ for power spectrum Eq.(28).
-Finally, we can calculate the $1\,\sigma$ error for a parameter $p_{\mu}$ as $\sigma(p_{\mu})=\sqrt{(\mathcal{F}^{-1})_{\mu\mu}}$, as shown in Table 2.
+Finally, we can calculate the $1\,\sigma$ marginalized error on a parameter $p_{\mu}$ as $\sigma(p_{\mu})=\sqrt{(\mathcal{F}^{-1})_{\mu\mu}}$, as shown in Table 2.
 III results
 Figure 1: Cluster number counts as a function of redshift $z$.
-Table 2: Forecasted $1\,\sigma$ errors for all parameters described in subsection II.5.
+Table 2: Marginalized $1\,\sigma$ errors forecasted for all the parameters described in subsection II.5.
 When a parameter is kept fixed, we use ’–’.
-The header labels the different cosmological models ($\nu\Lambda$CDM and $\nu w_{0}w_{a}$), scenarios (self-calibration and known SR), and probes (cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$).
-For complete descriptions, please refer to the main text.
+The header specifies the different cosmological models ($\nu\Lambda$CDM and $\nu w_{0}w_{a}$), scenarios for our knowledge of the scaling relation (self-calibration and known SR), and probes (cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$).There are almost no constraints from $P$ on the SR parameters, denoted by ’/’, and adding priors to these parameters does not affect the main results.
+For complete description, please refer to the main text.
 model $\nu\Lambda$CDM $\nu w_{0}w_{a}$
@@
 $\sigma_{8}$ 0.258 0.018 0.010 0.010 0.004 0.002 0.269 0.036 0.010 0.058 0.006 0.003
-$A_{\lambda}$ 1.5e+02 1.1e+07 1.1e+01 - - - 2.0e+02 1.1e+07 1.2e+01 - - -
-$B_{\lambda}$ 1.024 1.7e+05 0.203 - - - 1.356 1.7e+05 0.207 - - -
-$C_{\lambda}$ 1.628 1.1e+05 0.328 - - - 2.294 1.0e+05 0.336 - - -
-$D_{\lambda}$ 0.363 7.1e+04 0.074 - - - 0.415 7.0e+04 0.075 - - -
-$\kappa$ 0.466 5.4e+03 0.129 - - - 0.539 5.3e+03 0.136 - - -
-$B_{M0}$ 0.860 6.4e+01 0.029 - - - 1.006 6.3e+01 0.033 - - -
-$\alpha$ 0.431 1.9e+03 0.045 - - - 0.434 1.9e+03 0.047 - - -
+$A_{\lambda}$ 1.5e+02 / 1.1e+01 - - - 2.0e+02 / 1.2e+01 - - -
+$B_{\lambda}$ 1.024 / 0.203 - - - 1.356 / 0.207 - - -
+$C_{\lambda}$ 1.628 / 0.328 - - - 2.294 / 0.336 - - -
+$D_{\lambda}$ 0.363 / 0.074 - - - 0.415 / 0.075 - - -
+$\kappa$ 0.466 / 0.129 - - - 0.539 / 0.136 - - -
+$B_{M0}$ 0.860 / 0.029 - - - 1.006 / 0.033 - - -
+$\alpha$ 0.431 / 0.045 - - - 0.434 / 0.047 - - -
 $M_{\nu}(eV)$ 5.215 0.201 0.117 0.706 0.194 0.031 5.458 0.293 0.127 1.095 0.223 0.034
-In this section, we present the main results based on the CSST survey.
-The redshift distribution of CSST clusters is shown in Figure 1, with a total number count of 408,037.
-The predicted $1\,\sigma$ errors for all parameters, especially for the total neutrino mass $M_{\nu}$, are shown in Table 2, for the various probes, cosmological models, and scenarios.
-Specifically, these constraints is derived from the cluster number counts $N$, the cluster power spectrum $P$, and their combination $N+P$, in the two cosmological models: the $\nu\Lambda$CDM model and the $\nu w_{0}w_{a}$ model.
-To explore the impact of the main systematics in cluster cosmology, i.e.
+In this section, we present the main results forecasted for the CSST.
+The redshift distribution of the CSST clusters is shown in Figure 1, with a total number count of 408,037.
+(This is slightly different from our previous prediction of 415,000 in [30] due to the slight difference in the fiducial parameters.) The predicted $1\,\sigma$ marginalized errors for all parameters, especially for the total neutrino mass $M_{\nu}$, are shown in Table 2, for different probes, cosmological models, and scenarios.
+Specifically, these constraints are derived from the cluster number counts $N$, the cluster power spectrum $P$, and their combination $N+P$, and for the two cosmological models: the $\nu\Lambda$CDM model and the $\nu w_{0}w_{a}$ model.
+To address the impact of the major systematic in cluster cosmology, i.e.
 the uncertainty in the SR, we consider two scenarios: (1) the self-calibration scenario, where the SR parameters are jointly constrained with the cosmological parameters; and (2) the known SR scenario, where the SR parameters are fixed.
-To express clearly, the constraints on $M_{\nu}$ obtained from $N$, $P$ and $N+P$ are denoted as $\sigma^{N}(M_{\nu})$, $\sigma^{P}(M_{\nu})$ and $\sigma^{N+P}(M_{\nu})$ respectively.
+For clarity, the constraints on $M_{\nu}$ obtained from $N$, $P$ and $N+P$ are denoted as $\sigma^{N}(M_{\nu})$, $\sigma^{P}(M_{\nu})$ and $\sigma^{N+P}(M_{\nu})$ respectively.
 III.1 $\nu\Lambda$CDM model
-Figure 2: Constraints on cosmological parameters in the $\nu\Lambda$CDM model, with marginalizing over the SR parameters if exit.
-All contours denote 68% C.L.
-Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$.
-Solid and dotted lines represent self-calibration and known SR scenario respectively.
-We begin with the $\nu\Lambda$CDM model, referring to columns 2-7 in Table 2.
+Figure 2: Marginalized constraints on cosmological parameters for the $\nu\Lambda$CDM model.
+All contours show the 68% confidence regions.
+Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$ in turn.
+Solid and dotted lines represent the self-calibration and known SR scenarios respectively.
+We begin with the $\nu\Lambda$CDM model, whose constraints are shown in columns 2-7 in Table 2.
 The corresponding contour plot is shown in Fig.
 2 for the cosmological parameters $\{M_{\nu},\Omega_{b}h^{2},\Omega_{m}h^{2},\Omega_{DE},n_{s},\sigma_{8}\}$.
-It is evident that constraints on all cosmological parameters primarily arise from $P$, except for $\Omega_{DE}$ in the known SR scenario.
-This is because $N$ is only sensitive to the integration of the matter power spectrum, rather than its shape, resulting in weak constraints on $\Omega_{b}h^{2}$, $n_{s}$ and $M_{\nu}$ [40, 21, 72].
-Consequently, the constraint on $M_{\nu}$ from $N$ primarily stems from the influence of neutrinos on the amplitude of the matter power spectrum, leading to a strong degeneracy between $M_{\nu}$ and $\sigma_{8}$.
-In summary, the constraint on $M_{\nu}$ from $N$ is very weak, only $\sigma^{N}(M_{\nu})=5.215\ (0.706)$ eV in the self-calibration (known SR) scenario.
-The power spectrum $P$ serves as a more sensitive probe of the neutrino mass compared to $N$, providing a constraint of $\sigma^{P}(M_{\nu})=0.201\ (0.706)$ eV on its own.
+It is evident that constraints on all cosmological parameters primarily come from $P$, except for $\Omega_{DE}$ in the known SR scenario.
+This is because $N$ is sensitive to the integration of the cdm+b power spectrum, rather than its shape, resulting in weaker constraints on $\Omega_{b}h^{2}$, $n_{s}$ and $M_{\nu}$ [40, 20, 76].
+Consequently, the constraint on $M_{\nu}$ from $N$ primarily stems from the influence of neutrinos on the amplitude of the cdm+b power spectrum, leading to a strong degeneracy between $M_{\nu}$ and $\sigma_{8}$.
+In summary, the constraint on $M_{\nu}$ from $N$ is very weak, $\sigma^{N}(M_{\nu})=5.215\ (0.706)$ eV in the self-calibration (known SR) scenario.
+The power spectrum $P$ proves to be a more sensitive probe of the neutrino mass compared to $N$, providing a constraint of $\sigma^{P}(M_{\nu})=0.201\ (0.194)$ eV on its own.
 When combined with $N$, the improvement factor in the self-calibration (known SR) scenario is 1.72 (6.26).
 The incorporation of SR information has a substantial impact on $N$, resulting in an enhancement of constraints on all cosmological parameters by a factor of at least four.
-Particularly, the constraint on $M_{\nu}$ experiences a significant strength by a factor of 7.4, shifting from $5.215$ to $0.706$ eV.
-In contrast, the influence of SR information on $P$ is minimal, which is reflected in large errors of SR parameters in the third column, since the SR parameters only affects $b_{\text{eff}}$.
-Consequently, the enhance in the constraint on $M_{\nu}$ is only 1.04 times, from $0.201$ to $0.194$ eV, when incorporating SR information.
-In the combined analysis, the inclusion of SR information strengthens the constraint from $0.117$ to $0.031$ eV, which corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
+Particularly, the constraint on $M_{\nu}$ experiences a significant improvement by a factor of 7.4, shifting from $5.215$ to $0.706$ eV.
+In contrast, the influence of SR information on $P$ is less significant, which is reflected in the large errors of SR parameters in the third column, since the SR parameters only affect $b_{\text{eff}}$.
+Consequently, the improvement in the constraint on $M_{\nu}$ is only a factor of 1.04, from $0.201$ to $0.194$ eV, when incorporating SR information.
+In the combined analysis, the inclusion of SR information strengthens the constraint from $0.117$ to $0.031$ eV, which corresponds to a $\sim 2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
 III.2 $\nu w_{0}w_{a}$ model
-Figure 3: Similar to Fig. 2, but in $\nu w_{0}w_{a}$ model.
-Next we extend the $\nu\Lambda$CDM model to the $\nu w_{0}w_{a}$ model in order to investigate the impact of the dynamic dark energy on the neutrino constraint, referring to columns 8-13 in Table 2.
+Figure 3: Similar to Fig. 2, but for the $\nu w_{0}w_{a}$ model.
+Next we extend the $\nu\Lambda$CDM model to the $\nu w_{0}w_{a}$ model in order to investigate the impact of dynamic dark energy on the neutrino constraint, results showing in columns 8-13 in Table 2.
 The corresponding contour plot is shown in Fig.
-3 for the cosmological parameters $\{M_{\nu},w_{0},w_{a}\}$ with other parameters marginalized over.
-It is expected that including the parameters $\{w_{0},w_{a}\}$ will loosen the constraint on $M_{\nu}$ since both dark energy and massive neutrinos influence the growth rate of structures in the Universe.
+3 for the cosmological parameters $\{M_{\nu},w_{0},w_{a}\}$ with the other parameters marginalized over.
+It is expected that including the parameters $\{w_{0},w_{a}\}$ will weaken the constraint on $M_{\nu}$ since both dark energy and massive neutrinos affect the growth rate of structures in the Universe.
 As depicted in Fig.
-3, there is a small positive correlation between $M_{\nu}$ and $w_{0}$, as well as a distinct negative correlation between $M_{\nu}$ and $w_{a}$, for both $N$ and $P$.
-However, the direction of degeneracy is slightly different between $N$ and $P$, resulting in the constraint from the combination $N+P$ being less affected by the presence of the dynamic dark energy [73].
+3, there is a positive correlation between $M_{\nu}$ and $w_{0}$, as well as a negative correlation between $M_{\nu}$ and $w_{a}$, for both $N$ and $P$.
+However, the direction of degeneracy is different between $N$ and $P$, resulting in the constraint on $M_{\nu}$ from the combination $N+P$ being less affected by the presence of the dynamic dark energy [77].
 Consider the known SR scenario as an example.
-Adding $\{w_{0},w_{a}\}$ results in a loosening of $\sigma(M_{\nu})$ by a factor of 1.552 and 1.1471 for $N$ and $P$, respectively.
-In contrast, the combination $N+P$ demonstrates a loosening factor of only 1.079.
-The same holds true for the self-calibration scenario, where the loosening factor for the combination $N+P$ is 1.084.
-In summary, the constraint on the total neutrino mass, $M_{\nu}$, is mainly contributed by the cluster power spectrum, $P$.
-The inclusion of dynamic dark energy has minimal impact, whereas the availability of perfect SR knowledge plays a significant role.
+Adding $\{w_{0},w_{a}\}$ results in a weakening of $\sigma(M_{\nu})$ by a factor of 1.55 and 1.15 for $N$ and $P$, respectively.
+In contrast, the combination $N+P$ shows a weakening factor of only 1.10.
+The same holds true for the self-calibration scenario, where the weakening factor for the combination $N+P$ is 1.08.
+In summary, the total neutrino mass, $M_{\nu}$, is much better constrained by the cluster power spectrum than by the cluster number counts.
+The two probes are highly complementary in constraining $M_{\nu}$.
+The inclusion of dynamic dark energy has negligible impact on the joint neutrino mass constraint from $N+P$.
+Availability of perfect SR knowledge plays a significant role in constraining $M_{\nu}$.
 IV discussions
-Table 3: Forecasted the $1\,\sigma$ marginalized error for the the total neutrino for different conditions.
-The 4th row represents the fiducial result, same as shown in Table 2.
-The subsequent rows demonstrate the results under specific conditions: partial SR knowledge (5th row), $z_{max}=2$ (6th row), $\sigma_{z}=0.01$ (7th row), HMF uncertainty (8th row), bias uncertainty (9th row), and without considering the GISDB effect (10th row).
-The symbol ’/’ denotes that the constraint is the same as the fiducial condition.
+Table 3: The $1\,\sigma$ marginalized errors forecasted for the the total neutrino mass under different conditions.
+The 4th row give the fiducial results, as shown in Table 2.
+The following rows show the results under the conditions of: partial SR knowledge (5th row), $z_{max}=2$ (6th row), $\sigma_{z}=0.01$ (7th row), HMF uncertainty (8th row), bias uncertainty (9th row), and without considering the GISDB effect (10th row).
+The symbol ‘/’ represents that the constraint is the same as the fiducial case.
 model $\nu\Lambda$CDM $\nu w_{0}w_{a}$
@@
 w/o GISDB / 0.187 0.140 / 0.158 0.064 / 0.277 0.191 / 0.229 0.075
-In this section, we analyze several systematic effects and survey strategies, as presented in Table 3, and conclude by comparing our forecast with results from other surveys at the end of the section.
+In this section, we study several effects that may change our forecasts, including knowledge on SR parameters, the maximum redshift, uncertainties in redshift measurements, HMF and bias, as well as the GISDB effect, with results presenting in Table 3, and compare with literature at the end of this section.
 IV.1 SR parameters
@@
 The upper and the lower panel represents the constraint from $N$ and $N+P$ respectively.
-In our analysis, we do not directly incorporate the mass observable in CSST, i.e.
-the optical richness, into our model.
+In our analysis, we do not directly use the mass observable for CSST, i.e.
+the optical richness.
 Although the mass-richness relation (MRR) has been calibrated in numerous works (e.g.
-[32, 33, 34, 35, 36]), it will be affected by different surveys and cluster finders [74, 75].
-Therefore it is inappropriate to select a MRR calibrated in one specific survey to another survey.
-Instead, we employ a simple parameterization to account for our current lack of knowledge regarding the details of the mass-observable relations and their potential dependence on redshift and mass.
-However, our approach is not entirely devoid of information.
-Rather than treating $\sigma_{\ln M}$ as a nuisance parameter, we relate it with $\sigma_{\ln\lambda}$, and assume $\sigma_{\ln\lambda}$ comprises a combination of a constant intrinsic scatter and a Poisson-like term.
-In our main analysis, we have selected a MRR to model the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$.
-However, in actual surveys, we will utilize a specific MRR to replace Eq.(10), and there will be no need for an additional MRR to account for the Poisson-like term $1/\exp\left<\ln\lambda^{ob}(M)\right>$.
-Consequently, we fix parameters $\{A_{\lambda},B_{\lambda},C_{\lambda}\}$ in Eq.(II.1), while retaining $B_{\lambda}$ as a nuisance parameter in Eq.(14) to mimic this scenario, which we refer to as the partial self-calibration scenario.
+[32, 33, 34, 35, 36]), it will be affected by survey specifics and cluster finders [78, 79].
+Therefore it is inappropriate to apply a MRR calibrated in one specific survey to another survey.
+Instead, we employ a simple parametrization to account for our current lack of knowledge about the details of the mass-observable relations and their potential dependences on redshift and mass.
+However, our approach is not completely devoid of information.
+Rather than treating $\sigma_{\ln M}$ as a nuisance parameter, we relate it with $\sigma_{\ln\lambda}$, and assume $\sigma_{\ln\lambda}^{2}$ to be composed of a constant intrinsic scatter term and a Poisson-like term.
+In our default analysis, we have selected a MRR to model the Poisson-like term of $1/\exp\left<\ln\lambda^{ob}(M)\right>$.
+However, in real surveys, we can also use the richnesses for the detected clusters directly, then there will be no need for an additional MRR to specify the Poisson-like term.
+Consequently, we try fixing the parameters $\{A_{\lambda},B_{\lambda},C_{\lambda}\}$ in Eq.(II.1) while retaining $B_{\lambda}$ as a free parameter in Eq.(14) to describe this scenario, which we refer to as the partial self-calibration scenario.
 The 5th row in Table 3 shows the result.
 Similar to the previous section, SR parameters have a greater impact on $N$ than on $P$.
-The partial self-calibration scenario improve the constraint on $M_{\nu}$ from $N$ by a factor of 1.8 (1.2) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model.
+In the partial self-calibration scenario, the constraint on $M_{\nu}$ is improved by a factor of 1.8 (1.2) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model for $N$.
 However, the improvement is minimal for the constraint from $P$ or $N+P$.
@@
 On the other hand, the lower panel demonstrates that $\sigma^{N+P}(M_{\nu})$ is more sensitive to the scatter parameters $B_{\lambda},D_{\lambda}$ and $\kappa$, which suggests we’d better select an observable with smaller scatter when using the combination probe $N+P$ to constrain the total neutrino mass.
-Recently, we calibrate the intrinsic mass-richness relation of clusters in THE THREE HUNDRED hydrodynamic simulations [76].
-Our findings indicate that employing a skewed Gaussian distribution to model the richness at a fixed halo mass results in a simpler and smaller scatter compared to the conventional use of a log-normal distribution.
-Once this scaling relation is validated with the CSST mock data, we intend to incorporate it into future cosmological analyses.
+Recently, we have calibrated the intrinsic mass-richness relation of clusters with THE THREE HUNDRED hydrodynamic simulations [80].
+Our findings indicate that employing a skewed Gaussian distribution to model the richness at a fixed halo mass results in a more accurate compared to the conventional use of a log-normal distribution.
+We plan to incorporate these findings in future cluster analyses.
 IV.2 $z_{max}$
@@
 In our fiducial analysis, we assume a maximum redshift of $z_{\max}=1.5$ for cluster detection with CSST.
-However, by improving the cluster selection algorithm or jointing analysis with other surveys like Euclid [31], it may be possible to extend the maximum redshift to $z_{\max}=2$ [30].
+However, by improving the cluster selection algorithm or combining data from other surveys like Euclid [31], it may be possible to extend $z_{\max}=1.5$ to be even higher such as to $z_{\max}=2$ [30].
 The 6th row in Table 3 shows the result with $z_{max}=2$.
-Compared to the fiducial result, the total number counts is increased to 434534, and the constraint from $N$ has been improved by a factor of $\sim$1.7 (1.1) in the self-calibration (known SR) scenario for both two cosmological models.
-However, the improvement for the constraint from $P$ is very limited, with a factor of only $\sim$1.02 across all cases.
-Since the constraint on $M_{\nu}$ primarily comes from $P$, the improvement in $\sigma^{N+P}(M_{\nu})$ is not significant.
-In the self-calibration (known SR) scenario, the improvement is $\sim$1.1 (1.03) for both cosmological models.
-We also investigate a wide range of $z_{max}$, as depicted in Fig.
+Compared to the fiducial result, the total number count is increased to 434534, and the constraint from $N$ has been improved by a factor of $\sim$1.7 (1.1) in the self-calibration (known SR) scenario for both cosmological models.
+However, the improvement for the constraint from $P$ is very limited, with a factor of only $\sim$1.02 in all cases, probably due to the large shot noises in the higher redshift bins.
+Since the constraint on $M_{\nu}$ is dominated by $P$, the improvement in $\sigma^{N+P}(M_{\nu})$ is not significant either.
+In the self-calibration (known SR) scenario, the improvement is only $\sim$1.1 (1.03) for both cosmological models.
+We also investigate a wide range of $z_{max}$, results showing in Fig.
 5.
-The upper panel shows the $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model.
-It is evident that $\sigma^{N}(M_{\nu})$ decreases significantly with the increase of $z_{max}$.
-Conversely, $\sigma^{P}(M_{\nu})$ remains relatively constant once $z_{max}$ reaches $\sim$1.4.
-This constancy arises due to the exponential decrease in the cluster number density $\bar{n}$ Eq.(31) with increasing redshift, resulting in a very small effective volume $V_{\text{eff}}$ Eq.(II.4) at high redshifts.
+The upper panel shows $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model.
+$\sigma^{N}(M_{\nu})$ in the self-calibration scenario is too large to be shown in this figure, but it decreases significantly with the increase of $z_{max}$.
+Conversely, $\sigma^{P}(M_{\nu})$ remains relatively a constant after $z_{max}$ has reached $\sim$1.4.
+This constancy arises from the sharp drop in the cluster number density $\bar{n}$ Eq.(31) with increasing redshift, resulting in a very small effective volume $V_{\text{eff}}$ Eq.(II.4) at high redshifts.
 Consequently, the constraint from the combination $N+P$, i.e.
-$\sigma^{N+P}(M_{\nu})$, exhibits a moderate decrease with increasing $z_{\text{max}}$.
+$\sigma^{N+P}(M_{\nu})$, exhibits a much more moderate decrease with increasing $z_{\text{max}}$.
 The lower panel shows the ratio of $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ to $\sigma(M_{\nu})$ in $\nu\Lambda$CDM model.
@@
 IV.3 Redshift uncertainty
-Figure 6: The impact of the redshift uncertainty $\sigma_{z}$ on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model.
-Different colors (red, blue and black) represent different probes: the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$.
-Solid and dotted lines represent self-calibration and known SR scenario respectively.
-The redshift uncertainty $\sigma_{z}$ is optimistically assumed to be 0.001 in our fiducial analysis.
-This assumption is based on the availability of spectroscopic redshifts for all clusters observed by the CSST.
-However, in cases where this assumption does not hold, we consider a more conservative estimate of 0.01 for $\sigma_{z}$.
-The constraint result using this conservative estimate is presented in the 7th row of Table 3.
-We also extend $\sigma_{z}$ up to 0.03 in Fig.
-5, which is the photometric redshift error for Euclid at high redshift [72].
-The redshift uncertainty $\sigma_{z}$ has a minimal impact on $\sigma^{N}(M_{\nu})$ due to the redshift bin size.
-Even for the largest redshift uncertainty $\sim$0.03, it is still smaller than the redshift bin $\Delta z=0.05$ we choose.
+Figure 6: The impact of redshift uncertainty parameter $\sigma_{z}$ on $\sigma(M_{\nu})$ for $\nu w_{0}w_{a}$ CDM model.
+Different colors represent different probes: red for the cluster number counts $N$, blue for cluster power spectrum $P$, and black for their combination $N+P$.
+Solid and dotted lines represent self-calibration and known SR scenarios respectively.
+The redshift uncertainty parameter $\sigma_{z}$ is optimistically assumed to be 0.001 in our fiducial analysis.
+This accuracy probably highly rely on the availability of accurate spectroscopic redshifts for all clusters observed by the CSST.
+In case it will be challenging to achieve such an accuracy, we consider a more conservative estimate of 0.01 for $\sigma_{z}$.
+The constraint using this conservative estimate is presented in the 7th row of Table 3.
+We also extend $\sigma_{z}$ up to 0.03 with results displayed in Fig.
+6 for the $\nu w_{0}w_{a}$ specifically.
+Similarly to Fig.
+5, $\sigma^{N}(M_{\nu})$ in the self-calibration scenario is too large to be displayed in this figure.
+For both scenarios, the redshift uncertainty $\sigma_{z}$ has a minimal impact on $\sigma^{N}(M_{\nu})$ due to the relatively large redshift bin size.
 However, the influence of $\sigma_{z}$ on $\sigma^{P}(M_{\nu})$ is noteworthy.
-This is attributed to the damping effect caused by the redshift uncertainty, which leads to the partial erasure of information in $P$.
-It is observed that the behavior of $\sigma^{N+P}(M_{\nu})$ aligns with that of $\sigma^{P}(M_{\nu})$.
+This is attributed to the damping effect caused by the redshift uncertainty, which suppresses $P$ on relatively smaller scales.
+The behavior of $\sigma^{N+P}(M_{\nu})$ with increasing $\sigma_{z}$ is similar to that of $\sigma^{P}(M_{\nu})$, but less significant.
 Quantitatively, the ratio between $\sigma^{N+P}(M_{\nu})$ with $\sigma_{z}=0.01$ and $\sigma_{z}=0.001$ is $\sim$2.
-These findings underscore the significance of precise redshift measurements in effectively constraining the total neutrino mass.
+These findings underscore the importance of precise redshift measurements in effectively constraining the total neutrino mass.
 IV.4 HMF uncertainty
-Although the CDM prescription we use in subsection II.3 has alleviate the impact of neutrinos on the universality of HMF, the HMF of Tinker we employ is accurate at the $\sim$5level [26].
-Additionally, baryonic effects introduce an uncertainty of $\sim$10to the HMF [78, 79].
+Although the CDM prescription we adopt in subsection II.3 has alleviated the impact of massive neutrinos on the universality of HMF, the HMF of Tinker et al (2008) we employ is accurate at the $\sim$5level [26].
+Additionally, systematics such as different halo finders, non-$\Lambda$CDM simulations, baryonic effects etc.
+introduce extra uncertainty to the HMF [44, 82, 83].
 To account for these uncertainties, we incorporate two additional parameters $\{s,q\}$ to characterize the deviation of the Tinker’s HMF from the true HMF, following [84] and [85]:
 $\frac{dn}{d\ln M}=\left(\frac{dn}{d\ln M}\right)_{\text{Tinker}}\left[s\log\left(\frac{M}{10^{13.8}{{\,h^{-1}{\rm Mpc}}}}\right)+q\right].$ (36)
-The fiducial values are $s=0$ and $q=1$, degenerating the HMF to the Tinker’s.
-The 8th row in Table 3 presents the constraints with additional HMF uncertainties.
-Taking the $\nu w_{0}w_{a}$ model as an example, compared to $\sigma(M_{\nu})$ in the fiducial condition, the constraints from $N$ have been relaxed by a factor of 1.2 (2.7) in the self-calibration (known SR) scenario, while the constraints from $P$ remain largely unchanged.
-However for the constraints from $N+P$, the relaxation becomes more significant by a factor of 1.85 (3.97).
-Similar conclusions hold for the $\nu\Lambda$CDM model.
+The fiducial values are $s=0$ and $q=1$, reducing to the Tinker’s mass function.
+The 8th row in Table 3 shows the constraints with additional HMF uncertainties.
+Taking the $\nu w_{0}w_{a}$ model as an example, compared to $\sigma(M_{\nu})$ in the fiducial case, the constraints from $N$ have been relaxed by a factor of 1.2 (2.7) in the self-calibration (known SR) scenario, while the constraints from $P$ remain largely unchanged.
+However for the constraints from $N+P$, the degradation becomes more significant by a factor of 1.85 (3.97).
+For the $\nu\Lambda$CDM model, $\sigma(M_{\nu})$ from $N+P$ degrades by a factor of 1.09(2.13).
 As a result, the accuracy of the HMF plays a crucial role in constraining $M_{\nu}$ using $N+P$.
-It is worth mentioning that current and future emulators, based on simulations with massive neutrinos, have the potential to determine the HMF with greater precision, bypassing the reliance on a universal formula [82].
-IV.5 bias uncertainty
-Galaxy clusters are biased tracers of the CDM field.
-Fortunately, compared to galaxies, we have theory to describe the bias $b(M,z)$ in Eq.(17) for clusters, and the formula can be calibrated through simulations, similar to the HMF.
-Therefore, the bias of Tinker that we utilize also exhibits uncertainties.
-Moreover, in real surveys, extra bias will be introduced by selection effects, primarily due to projection effects and orientation biases.
-We incorporate two parameters $\{b_{s0},b_{s1}\}$ to model these uncertainties, following [83, 84]:
+It is worth mentioning that emulators for HMF, based on simulations with massive neutrinos, have the potential to achieve greater precision, bypassing the reliance on a universal formula [86].
+IV.5 Bias systematics
+Galaxy clusters are biased tracers of matter field.
+Fortunately, compared to galaxies, we have theory to describe the bias for clusters, and the formula can be calibrated with simulations, similar to the HMF.
+The bias of Tinker et al (2010) that we utilize already exhibits uncertainties.
+Moreover, in real surveys, extra bias will be introduced by cluster selection effects.
+For example, the redMaPPer clusters show an additional mass-dependent bias due to projection effects and orientation biases[87].
+We introduce two parameters $\{b_{s0},b_{s1}\}$ to model systematics on cluster bias, following [87, 88]:
 $b(M,z)=b_{\text{Tinker}}(M,z)b_{s0}\left(\frac{M}{5\times 10^{14}{{\,h^{-1}{\rm{M_{\odot}}}}}}\right)^{b_{s1}},$ (37)
 with fiducial values $b_{s0}=1,b_{s1}=0$.
-The constraint on $M_{\nu}$ with the bias uncertainty is presented in the 9 th row in the Table 3.
-For $\sigma^{P}(M_{\nu})$, in the self-calibration scenario, the effect of bias uncertainty is overshadowed by the effect of SR parameters.
+The constraint on $M_{\nu}$ with the bias uncertainty is presented on the 9 th row of Table 3.
+For $\sigma^{P}(M_{\nu})$, in the self-calibration scenario, the effect of bias uncertainty is completely negligible.
 However, in the known SR scenario, the inclusion of bias uncertainty relaxes the constraint by a factor of $\sim$1.03 (1.3) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model.
@@
 In the known SR scenario, it is 2.94 (3.5) in the $\nu\Lambda$CDM ($\nu w_{0}w_{a}$) model.
-These findings highlight the greater influence of bias uncertainty compared to that of the HMF.
-It emphasizes the need for careful consideration of bias uncertainties in future analyses.
+These findings highlight the importance of bias systematics to cluster cosmology, and emphasize the need for careful study of bias systematics in future analyses.
 IV.6 GISDB
-Figure 7: Bias.
-The impact of the GISDB effect on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model.
+Figure 7: The impact of the GISDB effect on $\sigma(M_{\nu})$ in $\nu w_{0}w_{a}$ CDM model.
 Different colors (blue and black) represent different probes: cluster power spectrum $P$ and the combination $N+P$.
 Solid and hollow dots represent results with or without GISDB effect.
-Each column represents the self-calibration scenario (SC), the known SR scenario (SR), SC with the bias uncertainty and SR with the bias uncertainty, respectively.
-In this section, we investigate the impact of the GISDB effect, which has been omitted in previous studies (e.g. [72, 31, 85]). The constraint without GISDB is presented in the 10th row of Table 3.
-GISDB effect compensates for the suppression of the power spectrum caused by neutrinos at small scales.As a result, we anticipate that the constraint obtained from $P$ will be relaxed when GISDB is considered.
-However, there is an exception, $\sigma^{P}(M_{\nu})$ changing from 0.229 to 0.223.
-The reason for this exception is that when using $P$ to constrain $M_{\nu}$, we utilize information not only from the shape of the power spectrum but also from its amplitude, predominantly through the effective bias $b_{\text{eff}}$.
+From left to right, the columns represent the self-calibration scenario (SC), the known SR scenario (SR), SC with the bias uncertainty and SR with the bias uncertainty, respectively.
+In this subsection, we investigate the impact of the GISDB, which has been missed in previous cluster studies (e.g. [76, 31, 89]). The constraint without GISDB is presented in the 10th row of Table 3.
+GISDB compensates partly for the suppression of the power spectrum caused by massive neutrinos on small scales.
+As a result, we expect that the constraint obtained from $P$ will be degraded when GISDB is considered.
+This is indeed what we find except for the “known SR” case in the $\nu w_{0}w_{a}$ model, where $\sigma^{P}(M_{\nu})$ gets slightly better from 0.229 to 0.223.
+When using $P$ to constrain $M_{\nu}$, we utilize information not only from the shape of the power spectrum but also from its amplitude, predominantly through the effective bias $b_{\text{eff}}$.
 By adding bias uncertainties, as discussed in the previous subsection, the amplitude information is blurred, leaving only the shape information.
@@
 Especially for $\sigma^{N+P}(M_{\nu})$, considering GISDB will strengthen the constraint by a factor of $\sim$1.5 (2.2) in the self-calibration (known SR) scenario.
-IV.7 Comparison with other cluster surveys
-In this subsection, we compare our results for CSST with others for LSST, Euclid, and CMB-S4.
-All the constraints discussed here are derived in the $\nu w_{0}w_{a}$ model, except for LSST, which is in the $\nu w_{0}$ model.
+IV.7 Comparison with other forecasts
+In this subsection, we compare our results forecasted for the CSST with other works in the literature forecasted for e.g.
+the LSST, Euclid, and CMB-S4.
+All the constraints discussed here are derived for the $\nu w_{0}w_{a}$ model, except for LSST, which is for the $\nu w_{0}$ model.
 The first work using cluster number counts and cluster power spectrum to constrain the total neutrino mass was performed by Wang et al.
-for LSST, a shear-limited weak lensing survey with a sky area of 18,000 deg2 and a redshift range of $z=[0.1,1.4]$.
-They forecast $\sigma^{P}(M_{\nu})=0.71$ eV and $\sigma^{N+P}(M_{\nu})=0.42$ eV, with an improvement factor of $\sim$1.34 when adding $N$.
+(2005) for several cluster surveys including the LSST, which can provide a shear-selected cluster catalogue in a sky area of 18,000 deg2 and a redshift range of $z=[0.1,1.4]$.
+They forecast $\sigma^{P}(M_{\nu})=0.71$ eV and $\sigma^{N+P}(M_{\nu})=0.42$ eV, with an improvement factor of $\sim$1.70 when adding $N$.
 In the self-calibration scenario, we obtain $\sigma^{P}(M_{\nu})=0.293$ eV and $\sigma^{N+P}(M_{\nu})=0.127$ eV, with a improvement factor of $\sim$2.31.
-The main difference between our results and those of Wang et al.
-lies in the improvement factors of 1.34 and 2.31, respectively.
-This disparity can be attributed to the fact that the total cluster number counts forecasted by Wang et al.
-is approximately $2\times 10^{5}$, which is lower than our estimate of $4\times 10^{5}$.
-Next, we compare our results with those obtained by Sartoris et al.
-for Euclid, an optical survey with a sky area of 15,000 deg2 and a redshift range of $z=[0.2,2]$ [31].
-They use a selection threshold of $N_{500,c}/\sigma_{\text{field}}$, a ratio of $N_{500,c}$ the number of cluster galaxies contained within a sphere of certain radius and $\sigma_{\text{field}}$ the rms of the field galaxy counts.
+This difference can be attributed to the fact that the total cluster number counts forecasted by Wang et al.
+is approximately $2\times 10^{5}$, which is much lower than our estimate of $4\times 10^{5}$.
+Next, we compare our results with those obtained by Sartoris et al (2016) forecasted for the Euclid, which can provide an optical cluster catalogue with a sky area of 15,000 deg2 and a redshift range of $z=[0.2,2]$ [31].
+They use a selection threshold of $N_{500,c}/\sigma_{\text{field}}$, a ratio of $N_{500,c}$ the number of cluster galaxies contained within a certain radius to $\sigma_{\text{field}}$ the rms of the field galaxy counts.
 They forecast a total cluster number counts of $2\times 10^{6}$ and $2\times 10^{5}$ for $N_{500,c}/\sigma_{\text{field}}=3$ and 5 respectively.
-Our $4\times 10^{5}$ lies between them.
+Our cluster counts $4\times 10^{5}$ lies between these two numbers.
 For the former selection $N_{500,c}/\sigma_{\text{field}}=3$, they forecast $\sigma^{N+P}(M_{\nu})=0.140$ eV in the self-calibration scenario.
-When in the known SR scenario, it becomes 0.121 with an improvement factor of $\sim$1.16.
+In the known SR scenario, it becomes 0.121eV with an improvement factor of $\sim$1.16.
 Our main result is $\sigma^{N+P}(M_{\nu})=0.127$ eV in self-calibration, slightly smaller than $0.140$ eV.
-However, perfect knowledge of SR parameters bring our constraint greater improvement of $\sim$3.7 to $\sigma^{N+P}(M_{\nu})=0.034$ eV.
-When extending $z_{max}$ from 1.5 to 2, our results become smaller $\sigma^{N+P}(M_{\nu})=0.118\ (0.033)$ eV in self-calibration (known SR) scenario.
+However, perfect knowledge of SR parameters brings our constraint greater improvement by $\sim$3.7 to $\sigma^{N+P}(M_{\nu})=0.034$ eV.
+When extending $z_{max}$ from 1.5 to 2, our results become even smaller $\sigma^{N+P}(M_{\nu})=0.118\ (0.033)$ eV in self-calibration (known SR) scenario.
 It is worth noting that if we do not consider the GISDB effect as they do, our result becomes $\sigma^{N+P}(M_{\nu})=0.191\ (0.075)$ eV, larger than their $0.140$ eV.
-Finally, we compare our results with those obtained by Cromer et al.
-for CMB-S4, a Sunyaev-Zel’dovich (SZ) survey with a sky area of 10,000 deg2 and a redshift range of $z=[0.1,1.9]$ [85].
+Finally, we compare our results with forecasts by Cromer et al (2019) for the CMB-S4, which can provide a Sunyaev-Zel’dovich (SZ) cluster catalogue with a sky area of 10,000 deg2 and a redshift range of $z=[0.1,1.9]$ [89].
 They forecast a total cluster number counts of $\sim 10^{5}$.
-By combining Planck primary CMB and extra mass calibration from CMB lensing, they forecast $\sigma^{N+P}(M_{\nu})\sim 0.1$ eV, which falls between our results of $0.127$ eV and $0.034$ eV in the self-calibration and known SR scenarios, respectively.
+By combining with Planck primary CMB anisotropies and extra mass calibration from weak lensing, they obtain $\sigma^{N+P}(M_{\nu})\sim 0.1$ eV, which falls between our results of $0.127$ eV and $0.034$ eV in the self-calibration and known SR scenarios, respectively.
 It is worth noting that they find the improvement on $\sigma(M_{\nu})$ when adding $P$ is very small $\sim$0.5%.
-This may be due to the fact that they have already incorporated prior information from the primary CMB and additional mass calibration, resulting in the $N$ alone having a certain constraining ability.
-Furthermore, their redshift uncertainty is reported to be 0.01, which has a significant impact on the constraining ability of $P$.
-They also consider the bias uncertainty and find it has less than 1% effect on $\sigma(M_{\nu})$, mainly because $P$ has a minimal contribution on the constraint.
+This may be due to the fact that they have already incorporated information from the primary CMB and additional mass calibration in cluster number counts, making adding in $P$ less effective.
+Furthermore, their redshift uncertainty is set to be a conservative value of 0.01, which has a significant impact on the constraining ability of $P$.
+They also consider the bias uncertainty and find it has negligible effect on the combined constraint on $\sigma(M_{\nu})$, mainly because $P$ has minimal contribution to the constraint.
 V conclusions
-In this work, we forecast constraints on the total neutrino mass from CSST galaxy clusters.
-Specifically, we employ Fisher matrix analysis to derive the constraint $\sigma(M_{\nu})$ from the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$.
-In addition to considering the most basic cosmological extension, the standard cosmology plus neutrinos, $\nu\Lambda$CDM model, we also explore the inclusion of dynamic dark energy, $\nu w_{0}w_{a}$ model.
-Since the largest source of systematics in galaxy cluster cosmology arises from the mass-observable relation, we examine cases where the SR parameters are completely unknown and perfectly known as self-calibration scenario and known SR scenario.
-The main results are presented in Table 3.
-The results between $\nu\Lambda CDM$ and $\nu w_{0}w_{a}$ model are similar.
-Unless otherwise specified, the following results refer to $\nu w_{0}w_{a}$ model:
-  • CSST is expected to detect 408037 clusters with the mass threshold $M_{200m}\geq 0.836\times 10^{14}M_{\odot}/h$ and a survey area of 17,500 deg2 in the redshift range of $z=[0,1.5]$ , as depicted in Fig.
+In this work, we forecast constraints on the total neutrino mass from the CSST galaxy clusters.
+Specifically, we employ the Fisher matrix formalism to derive the constraint from the cluster number counts $N$, cluster power spectrum $P$, and their combination $N+P$.
+In addition to considering the simplest extension to the $\Lambda$CDM model, the $\nu\Lambda$CDM model ($\Lambda$CDM plus massive neutrinos), we also explore the inclusion of dynamic dark energy, the $\nu w_{0}w_{a}$ model.
+We carefully take into account the effects of massive neutrinos on the halo mass function and halo bias.
+Since the largest source of systematics in galaxy cluster cosmology arises from the mass-observable relation, we examine two cases where the SR parameters are completely unknown or perfectly known, named as the self-calibration scenario and known SR scenario respectively.
+Our main results are presented in Table 2 and Table 3.
+The results for the $\nu\Lambda CDM$ and $\nu w_{0}w_{a}$ model are similar.
+Unless otherwise specified, the following results refer to the $\nu w_{0}w_{a}$ model:
+  • CSST is expected to detect $\sim 400,000$ clusters with a mass threshold of $M_{200m}=0.836\times 10^{14}M_{\odot}/h$ in a survey area of 17,500 deg2 and redshift range of $z=[0,1.5]$ , as depicted in Fig.
 1.
-  • Probe $N$ provides a weak constraint on $M_{\nu}$, $\sigma^{N}(M_{\nu})=5.458\ (1.095)$ eV in the self-calibration (known SR) scenario.
-The constraint is mainly contributed by $P$, with $\sigma^{P}(M_{\nu})=0.293\ (0.223)$ eV.
+  • $N$ provides a weak constraint on $M_{\nu}$, $\sigma^{N}(M_{\nu})=5.458\ (1.095)$ eV in the self-calibration (known SR) scenario.
+The constraint is dominated by $P$, with $\sigma^{P}(M_{\nu})=0.293\ (0.223)$ eV.
 However, when $N$ is added, the constraint can be improved by a factor of 2.31 (6.56), resulting in $\sigma^{N+P}(M_{\nu})=$ 0.127 (0.034) eV.
-  • Having knowledge of the SR parameters has a significant impact on $\sigma^{N}(M_{\nu})$ and consequently affects $\sigma^{N+P}(M_{\nu})$.
+  • Knowledge of the SR parameters has a significant impact on $\sigma^{N}(M_{\nu})$ and consequently affects $\sigma^{N+P}(M_{\nu})$.
 While partial knowledge leads to a small improvement in $\sigma^{N+P}(M_{\nu})$ from $0.127$ eV to $0.125$ eV, having perfect knowledge results in a substantial improvement to $0.034$ eV.
-This corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
-  • Increasing $z_{max}$ from 1.5 to 2.0 will detect more clusters, increasing the count from $4\times 10^{5}$ to $4.3\times 10^{5}$.
-While this change has a very small impact on $\sigma^{P}(M_{\nu})$, improving it by a factor of 1.02 (1.02) in the self-calibration (known SR) scenario, it significantly strengthens $\sigma^{N}(M_{\nu})$ by a factor of 1.64 (1.20).
-As a result, it also strengthens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.08 (1.03), leading to $\sigma^{N+P}(M_{\nu})=0.118(0.033)$ eV.
-  • Among the uncertainties we considered, the redshift uncertainty $\sigma_{z}$ has the largest impact on $\sigma^{P}(M_{\nu})$, relaxing the constraint by a factor of 2.23 (2.13) when increasing $\sigma_{z}$ from 0.001 to 0.01.
-Consequently, this leads to a relaxation in $\sigma^{N+P}(M_{\nu})$ by a factor of 1.90 (2.15) in the self-calibration (known SR) scenario.
-  • We also examine the impact of uncertainties in the HMF and bias.
-The HMF uncertainties primarily affect $\sigma^{N}(M_{\nu})$, while the bias uncertainties mainly affect $\sigma^{P}(M_{\nu})$.
-However, the effects of both uncertainties on $\sigma^{N+P}(M_{\nu})$ are comparable, leading to a relaxation of the constraints by a factor of approximately 2 (4) in the self-calibration (known SR) scenario
-  • Finally, we investigate the impact of the GISDB effect, which has been omitted in the previous work.
-This effect compensates for the suppression of the power spectrum caused by neutrinos on small scales, thereby relaxing the constraint on neutrinos.
-However, taking into account the degeneracy between $M_{\nu}$ and bias, as well as the addition of the probe $N$, our results demonstrate that considering this effect actually tightens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.50 (2.21) in the self-calibration (known SR) scenario.
-In summary, CSST clusters are expected to provide a constraint on $M_{\nu}$ of $0.127$ eV.
-With perfect knowledge of the SR parameters, this constraint can be strengthened to $0.034$ eV, enabling a $2\sigma$ detection of the minimum neutrino mass of 0.06 eV.
+This corresponds to a $\sim 2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
+  • Increasing $z_{max}$ from 1.5 to 2.0 will result in more detected clusters, increasing the total count from $4\times 10^{5}$ to $4.3\times 10^{5}$.
+While this change has a very small impact on $\sigma^{P}(M_{\nu})$, improving it by a factor of 1.02 (1.02) in the self-calibration (known SR) scenario, it does strengthen $\sigma^{N}(M_{\nu})$ by a factor of 1.64 (1.20).
+As a result, it strengthens $\sigma^{N+P}(M_{\nu})$ a little bit, by a factor of 1.08 (1.03), leading to $\sigma^{N+P}(M_{\nu})=0.118(0.033)$ eV.
+  • Among the systematics we considered, the redshift uncertainty $\sigma_{z}$ has the largest impact on $\sigma^{P}(M_{\nu})$, relaxing the constraint by a factor of 2.23 (2.13) when increasing $\sigma_{z}$ from 0.001 to 0.01.
+Consequently, this leads to a degradation in $\sigma^{N+P}(M_{\nu})$ by a factor of 1.90 (2.15) in the self-calibration (known SR) scenario.
+  • We also examine the impact of uncertainties in the HMF and cluster bias.
+The HMF uncertainties primarily affect $\sigma^{N}(M_{\nu})$, while the bias uncertainties only affect $\sigma^{P}(M_{\nu})$.
+However, the effects of both uncertainties on $\sigma^{N+P}(M_{\nu})$ are comparable, leading to a degradation of the constraints by a factor of approximately 2 (4) in the self-calibration (known SR) scenario
+  • Finally, we investigate the impact of the GISDB, which has been missed in previous studies.
+This effect compensates for the suppression of the power spectrum caused by massive neutrinos on small scales, thereby one would expect it relaxes the constraint on neutrino mass.
+However, taking into account the degeneracy between $M_{\nu}$ and bias, as well as the addition of the probe $N$, our results show that considering this effect actually tightens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.50 (2.21) in the self-calibration (known SR) scenario.
+In summary, we find CSST clusters can provide a constraint on $M_{\nu}$ of $\sim 0.1$ eV by combining N and P.
+With perfect knowledge of the SR parameters, this constraint can be tightened to $\sim 0.03$ eV, enabling a $2\sigma$ detection of the minimum neutrino mass of 0.06 eV.
 Recently, Lin et al.
@@
 By combining clusters and galaxies, it is expected that CSST will significantly strengthen the constraint on the total neutrino mass.
-Moreover, the emergence of new probes such as voids [87] and Minkowski functions [88, 89] has opened up additional avenues for studying neutrino masses.
-By combining these innovative probes with traditional methods, it is anticipated that in the near future it will be possible to accurately constrain neutrino masses and distinguish between different mass hierarchies.
+Moreover, the development of new probes such as voids [91] and Minkowski functionals [92, 93] has opened up new avenues for constraining neutrino masses.
+By combining these novel probes with traditional methods, it is expected that in the near future it will be possible to accurately constrain neutrino masses and distinguish between different mass hierarchies.
 Acknowledgements.
-This work is supported by the National Natural Science Foundation of China Grants No.
-12173036, 11773024, 11653002, 11421303, by the National Key R&\&&D Program of China Grant No.
-2021YFC2203100, by the China Manned Space Project Grant No.
+We thank Yan Gong for useful conversations.
+This work is supported by the National Key R&D Program of China Grant No.
+2022YFF0503404 and No.
+2021YFC2203100, by the National Natural Science Foundation of China Grants No.
+12173036 and 11773024, by the China Manned Space Project “Probing dark energy, modified gravity and cosmic structure formation by CSST multi-cosmological measurements” and Grant No.
 CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No.
-WK3440000004 and WK3440000005, and by the CAS Interdisciplinary Innovation Team.
+WK3440000004 and WK3440000005, by Cyrus Chun Ying Tang Foundations, and by the 111 Project for ”Observational and Theoretical Research on Dark Matter and Dark Energy” (B23042).